### v1

# Stopping rules for Monte Carlo methods of martingale difference type

Abstract
We establish a practical and easy-to-implement sequential stopping rule for the martingale central limit theorem, focusing on Monte Carlo methods for estimating the mean of a non-iid sequence of martingale difference type. Starting with an impractical scheme based on the standard martingale central limit theorem, we progressively address its limitations from implementation perspectives in the non-asymptotic regime. Along the way, we compare the proposed schemes with their counterparts in the asymptotic regime. The developed framework has potential applications in various domains, including stochastic gradient descent methods. Numerical results are provided to demonstrate the effectiveness of the developed stopping rules in terms of reliability and complexity.
Keywords: Monte Carlo methods; martingale central limit theorem; martingale difference sequences; sequential stopping rules; stochastic gradient descent; early stopping.
2020 Mathematics Subject Classification: 65C05, 60G40, 60G42, 62L15, 68T01.
1 Introduction
In Monte Carlo methods in general terms, it is a common practice to fix the total number of samples before running the simulation, while, without sufficient knowledge of the underlying distribution, it could be risky to terminate the procedure at such a priori fixed timing outright for the reason that the estimator that one is dealing with may converge extremely slowly. Moreover, the length of the simulation needs to be decided in reference to various budgetary constraints, as computing resources available are finite in practice without an exception. Those issues can be addressed by equipping the simulation procedure with a sequential (or, equivalently, adaptive) stopping rule, so that the total number of samples can be decided along the way depending on the previous computation, rather than stopping the simulation too soon or continuing the simulation endlessly.
In the framework of the standard iid Monte Carlo simulation based on the central limit theorem, there has been a great deal of interest in such stopping criteria in the literature for a very long time. In the asymptotic regime, sequential stopping rules were first developed based on the first and second moments and investigated in terms of consistency and convergence rates (for instance, [3, 12, 26, 29], to name just a few). Nevertheless, in the non-asymptotic regime where the cost of computing needs to be taken into account, such asymptotics-based stopping rules may not function as theoretically expected, without extra information and restrictions, for instance, higher moments and boundedness of the underlying distribution (see, for example, [1, 9, 17] and references therein), due to an inaccurate estimate of the error probability at an early stage of the simulation procedure.
Departing from the standard iid Monte Carlo simulation and its underlying standard central limit theorem, non-iid sequences have also attracted attention in the context of sequential stopping rules (for instance, [9, 10, 23]) and, more recently, machine learning. Among a variety of such non-iid sequences, the focus of the present work is set on a sequence of martingale difference type, which comes with the martingale central limit theorem. To be exact, we aim to establish a rigorous yet easy-to-implement sequential stopping rule for the martingale central limit theorem with a view towards Monte Carlo methods in estimating the mean of a sequence of random variables (just like iid Monte Carlo methods), whose law, however, keeps changing in reference to the past information within a martingale framework (unlike iid Monte Carlo methods).
The relevant situation arises, for instance, from adaptive variance reduction techniques employed concurrently with the computation of the empirical mean so that the underlying randomness can be used in both procedures, such as importance sampling [21, 22], control variates and stratified sampling [32], to mention just a few. Due to their remarkable effectiveness in reducing the estimator variance without requiring much additional computing effort for running the parallel procedure, those adaptive variance reduction methods have found various fields of application, such as structural engineering [30, 31] and financial risk management [11]. The proposed framework on stopping rules for martingale difference sequences also has significant potential in machine learning optimization, particularly stochastic gradient descent methods where the stochastic gradient noises naturally form a martingale difference sequence. While conventional stopping methods often rely on validation set performance [28], our approach provides a principled alternative based on the statistical properties of the optimization trajectory itself. This is particularly of relevance given the recent variance reduction techniques for stochastic gradient descent methods [18], complementing the existing developments on validation-free stopping criteria [5].
In constructing valid stopping rules of this martingale type, the primary challenge lies in the impracticality of applying the martingale central limit theorem in its standard form. Unlike the standard central limit theorem for iid Monte Carlo methods, where the limiting variance is merely unknown, the theoretical variance underpinning the martingale central limit theorem (such as in [25]) is not only unknown but also cumulative averaging. In addition, the theoretical variance here is unconditional, that is, poorly aligned with the conditional expectations inherent to martingales. Moreover, the martingale central limit theorem generally hinges on intricate and often unverifiable technical conditions, which further complicates its practical use. Given those difficulties, we explore an implementable stopping rule by progressively replacing the theoretical variance with more implementable alternatives, while carefully imposing technical conditions along the way for the corresponding stopping rules to remain valid. In particular, in its final form, the developed stopping rule operates upon a fully implementable empirical variance.
The rest of this paper is set out as follows. To begin with, in Section 2, we prepare the general notation and set up standing assumptions along with some relevant examples in our context. In Section 3, we describe the problem formulation and review the martingale central limit theorem in its standard form. We construct in Section 4 sequential stopping rules based on the martingale central limit theorem with its theoretical batch variance replaced with more implementable ones step by step, first with the conditional variance (Section 4.2.1) and then the empirical variance (Section 4.2.2). We provide numerical results in Section 5 to demonstrate the effectiveness of the proposed stopping rule on a time series (Section 5.1) and an adaptive variance reduction method (Section 5.2). We conclude the present study by highlighting future research directions in Section 6. To maintain the flow of the paper, we collect the proofs in Section 7. Finally, in the supplementary material, we provide verifications of the standing assumptions and conditions for the numerical examples examined in Section 5 and discuss relevant stopping rules to put the proposed framework into perspective.
2 Preliminaries
[…]
We reserve the notation $\{X_{k}\}_{k\in\mathbb{N}}$ for a sequence of random variables of interest. Based upon the sequence $\{X_{k}\}_{k\in\mathbb{N}}$, we work on a filtered probability space, denoted by $(\Omega,\mathcal{F},(\mathcal{F}_{k})_{k\in\mathbb{N}_{0}},\mathbb{P})$ on which a sequence of random variables $\{X_{k}\}_{k\in\mathbb{N}}$ is defined. In particular, the filtration starts from the $\sigma$-field $\mathcal{F}_{0}$, which is a collection of $\mathbb{P}$-null sets, that is $\mathcal{F}_{0}:=\sigma(\mathcal{N})$ with $\mathcal{N}:=\{N\subseteq\Omega:\exists A\in\vee_{k\in\mathbb{N}}\mathcal{F}_{k},\mathbb{P}(A)=0,N\subseteq A\}$. Note that the filtration $(\mathcal{F}_{k})_{k\in\mathbb{N}_{0}}$ is not necessarily generated by the sequence $\{X_{k}\}_{k\in\mathbb{N}}$. Moreover, for the sake of simplicity and convenience, we write $\mathbb{P}_{k}(\cdot):=\mathbb{P}(\cdot|\mathcal{F}_{k})$, $\mathbb{E}_{k}[\cdot]:=\mathbb{E}[\cdot|\mathcal{F}_{k}]$, and $\mathrm{Var}_{k}(\cdot):=\mathrm{Var}(\cdot|\mathcal{F}_{k})$ for all $k\in\mathbb{N}_{0}$. In addition, we let $\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}$ denote the convergence in probability under $\mathbb{P}_{0}$ and indicate by $G_{k}\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}+\infty$ (for a sequence of random variables $\{G_{k}\}_{k\in\mathbb{N}}$) that $\lim_{k\to+\infty}\mathbb{P}_{0}(|G_{k}|>c)=1$ for all $c>0$.
Now, we impose the base assumption on the sequence $\{X_{k}\}_{k\in\mathbb{N}}$ of random variables, which remains true throughout the present study.
Assumption 2.1.
For every $k\in\mathbb{N}$, the random variable $X_{k}$ is $\mathcal{F}_{k}$-measurable with
$\mathbb{E}_{k-1}\left[X_{k}\right]=\mu,$ (2.1)
for some $\mu\in\mathbb{R}$.
Clearly, the standing assumption (2.1) is much weaker than the usual iid assumption and not to indicate that the sequence $\{X_{k}\}_{k\in\mathbb{N}}$ is a martingale (for which it would have to satisfy $\mathbb{E}_{k-1}[X_{k}]=X_{k-1}$ for all $k\in\mathbb{N}$ with suitable $X_{0}\in\mathcal{F}_{0}$). Instead, the assumption (2.1) can be interpreted to form a wider class than martingale difference sequences since the sequence $\{X_{k}\}_{k\in\mathbb{N}}$ is a martingale difference sequence if, moreover, $\mu=0$. It is, however, inappropriate to assume $\mu=0$ upfront or consider $Y_{k}:=X_{k}-\mu$ instead, since the constant $\mu$ is deemed unknown and is exactly what one looks to estimate in the present context. We also note that a similar assumption has been made in [9, Assumption 1.1] in the context of relative-precision stopping rules, but along with a boundedness condition on the underlying randomness (that is, $X_{n}\in[0,1]$, $a.s.$).
To further illustrate the scope of the standing assumption (2.1), we describe a few relevant problem settings below.
Example 2.2.
(a) If $\{X_{k}\}_{k\in\mathbb{N}}$ is a sequence of iid random variables with finite mean $\mu$ and if $(\mathcal{F}_{k})_{k\in\mathbb{N}_{0}}$ is its natural filtration, then it satisfies Assumption 2.1. In this sense, our developments in what follows can be thought of as generalizations of the existing stopping rules for iid Monte Carlo methods. At this stage, however, we wish to emphasize that such sequences of iid random variables are outside the scope of the present study for the reason that such a way too simple situation can be treated based on much lighter conditions (for instance, [1, 3, 10, 17, 26, 29]) than we impose in what follows for the general problem setting (2.1).
(b) Consider an ARCH(1) sequence $\{X_{k}\}_{k\in\mathbb{N}}$, defined by $X_{k}:=(\beta+\alpha X_{k-1}^{2})^{1/2}V_{k}$ for some $\alpha\in(0,1)$ and $\beta\in(0,+\infty)$, where $\{V_{k}\}_{k\in\mathbb{N}}$ is a sequence of iid random variables with zero mean and unit variance. If $(\mathcal{F}_{k})_{k\in\mathbb{N}_{0}}$ is the natural filtration generated by the sequence $\{V_{k}\}_{k\in\mathbb{N}}$, then it holds true that $\mathbb{E}_{k-1}[X_{k}]=0$ for all $k\in\mathbb{N}$, which satisfies Assumption 2.1 with $\mu=0$. Later in Section 5.1, we examine our stopping rules on this problem setting.
(c) A variety of adaptive variance reduction methods (such as [21, 22]) are also within our scope, as described in the introductory section. Those frameworks can be formulated in common as follows. Let $\{U_{k}\}_{k\in\mathbb{N}}$ be a sequence of iid uniform random vectors on $(0,1)^{d}$ and let $(\mathcal{F}_{k})_{k\in\mathbb{N}_{0}}$ be its natural filtration. For all $k\in\mathbb{N}$, we set $X_{k}=R(U_{k};{\bm{\theta}}_{k-1})$, where $R(\cdot;{\bm{\theta}})$ is an unbiased estimator for $\mu$, in the sense of $\mathbb{E}_{0}[R(U;{\bm{\theta}})]=\mu$ for $U\sim U(0,1)^{d}$, irrespective of the parameter $\bm{\theta}$}, corresponding to importance sampling and/or control variates. If the sequence $\{{\bm{\theta}}_{k}\}_{k\in\mathbb{N}_{0}}$ of the parameters is constructed in an adaptive manner to the filtration $(\mathcal{F}_{k})_{k\in\mathbb{N}}$, then it holds true that $\mathbb{E}_{k-1}[X_{k}]=\mu$, due to ${\bm{\theta}}_{k-1}\in\mathcal{F}_{k-1}$ and $U_{k}\perp\mathcal{F}_{k-1}$ for all $k\in\mathbb{N}$. The stopping rules that we develop are examined on adaptive control variates in Section 5.2. ∎
Hereafter, we reserve the notation $t$ for the order of batches taking its value in $\mathbb{N}_{0}$. Fix a sequence of non-negative deterministic integers $\{m(t)\}_{t\in\mathbb{N}_{0}}$ satisfying $m(0)=0$ and m(t)>m(t1)m(t)>m(t-1) for all $t\in\mathbb{N}$, that is, $\lim_{t\to+\infty}m(t)=+\infty$. We denote by $M(t):=\{m(t-1)+1,\cdots,m(t)\}$ and its cardinality |M(t)|:=m(t)m(t1)|M(t)|:=m(t)-m(t-1), respectively, the set of indices in the $t$th batch and the $t$th batch size, for all $t\in\mathbb{N}$. We formalize the assumption that the size of batches $|M(t)|$ grows to infinity, as below.
Assumption 2.3.
We henceforth assume $|M(t)|\to+\infty$, as $t\to+\infty$.
Now, in order to describe where stopping rules may play a role, we first define the empirical batch mean by
$\mu(t):=\frac{1}{|M(t)|}\sum_{k\in M(t)}X_{k},\quad t\in\mathbb{N}.$ (2.2)
Henceforth, we indicate the batch index $t$ as a variable of function (that is, $(t)$), whether deterministic (such as $M(t)$ and $m(t)$) or probabilistic (such as $\mu(t)$), so as to distinguish from the subscript index $k$ (like, $X_{k}$).
Below, we state the standard results of (i) unbiasedness and (ii) strong consistency of the empirical batch mean (2.2), with a sketchy proof given in Section 7. Note that the denominator in the condition of (ii) is not $k^{2}$ but $(k-m(t-1))^{2}$, due to the batch arrangement $M(t)=\{m(t-1)+1,\cdots,m(t)\}$.
Theorem 2.4.
(i) It holds $\mathbb{P}_{0}$-$a.s.$ that $\mathbb{E}_{m(t-1)}[\mu(t)]=\mathbb{E}_{0}[\mu(t)]=\mu$ for all $t\in\mathbb{N}$.
(ii) If $\limsup_{t\to+\infty}\sum_{k\in M(t)}(k-m(t-1))^{-2}{\rm Var}_{0}(X_{k})<+\infty$, then it holds $\mathbb{P}_{0}$-$a.s.$ that $\mu(t)\to\mu$, as $t\to+\infty$.
The experiment (2.2) can be significantly different for different batches, because it is not based on iid realizations but on a single martingale difference sequence (not even an array), depending largely on its past history. Even so, Theorem 2.4 provides a firm theoretical base that ensures our problem setting makes perfect sense in estimating the unknown mean $\mu$, regardless of the batch in which one is computing the empirical batch mean (2.2).
3 Problem formulation
In light of the unbiasedness and strong consistency of the empirical batch mean (2.2) (Theorem 2.4), we aim to construct a suitable rule for stopping the simulation at the $t$th batch, in which the error probability $\mathbb{P}_{0}(|\mu(t)-\mu|>\epsilon)$ is satisfactorily small (than a preset acceptable error probability $\delta\in(0,1)$ with a preset precision level $\epsilon\in(0,+\infty)$) for the first time. Intuitively, we need to correspond smaller values of the precision $\epsilon$ and error probability $\delta$ to larger values of the stopping batch $\tau$. From a practical point of view, however, this criterion is not quite useful per se, as the error probability $\mathbb{P}_{0}(|\mu(t)-\mu|>\epsilon)$ is generally not assessable unless the law of the distance $|\mu(t)-\mu|$ is known. It is rather absurd to assume such knowledge, as one is conducting experiments because the constant $\mu$ is unknown in the first place.
On top of the non-assessability of the error probability, it is also a fatal drawback that the unbiasedness (Theorem 2.4 (i)) is generally no longer valid for a general $(\mathcal{F}_{m(t)})_{t\in\mathbb{N}_{0}}$-stopping time $\tau$, if the empirical batch mean is given as an output at the stopping batch (that is, $\mathbb{E}_{0}[\mu(\tau)]\neq\mu$). As an unrealistic yet illustrative example, suppose the support of the random variable $X_{k}$ under $\mathbb{P}_{0}$ is $\mathbb{R}$ for every $k\in\mathbb{N}$ and the stopping batch $\tau$ is defined as the first batch $t$ in which all elements $\{X_{k}\}_{k\in M(t)}$ are strictly positive. Then, for every $k\in M(\tau)$ in the stopping batch, the support of $X_{k}$ is $(0,+\infty)$. If $\mu=0$ (despite it is unknown in reality), then the original empirical batch mean at the stopping batch, which is now strictly positive by definition, is evidently biased for the constant $\mu$ of interest, that is, $\mathbb{E}_{0}[\mu(\tau)]>0=\mu$, $\mathbb{P}_{0}$-$a.s.$
[…]
$(X_{1},\cdots,X_{m(1)},X_{m(1)+1},\cdots,X_{m(\tau-1)},X^{\star}_{m(\tau-1)+1},\cdots,X^{\star}_{m(\tau)}),$
with which we compute and output the empirical batch mean $\mu_{\star}(\tau)$ of the resampled batch $(X^{\star}_{m(\tau-1)+1},\cdots,X^{\star}_{m(\tau)})$, defined by
$\mu_{\star}(\tau):=\frac{1}{|M(\tau)|}\sum_{k\in M(\tau)}X^{\star}_{k}.$ (3.1)
To distinguish the two random sources $\{X_{k}\}_{k\in M(\tau)}$ and $\{X^{\star}_{k}\}_{k\in M(\tau)}$ correctly, we hereafter denote by $\mathbb{P}^{\star}_{k-1}$ the probability measure on the $\sigma$-field $\mathcal{F}^{\star}_{k-1}$ that makes the branched sequence $(X_{1},\cdots,X_{m(\tau-1)},X^{\star}_{m(\tau-1)+1},\cdots,X^{\star}_{k-1})$ measurable for all $k\in M(\tau)$ and leaves the original stopping batch $(X_{m(\tau-1)+1},\cdots,X_{m(\tau)})$ out of its measurability. Accordingly, we write $\mathbb{E}_{k-1}^{\star}$ and ${\rm Var}_{k-1}^{\star}$, respectively, for the expectation and variance taken under the probability measure $\mathbb{P}^{\star}_{k-1}$. We note that $\mathcal{F}_{k-1}^{\star}=\mathcal{F}_{k-1}$, $\mathbb{P}_{k-1}^{\star}=\mathbb{P}_{k-1}$, $\mathbb{E}_{k-1}^{\star}=\mathbb{E}_{k-1}$ and ${\rm Var}_{k-1}^{\star}={\rm Var}_{k-1}$ if $k=m(\tau-1)+1$.
Prior to further analysis, we provide Figure 1 to illustrate the procedure for constructing the resampled empirical batch mean (3.1) along with examples below in Example 3.1.
[…]
Figure 1: Typical arrangement of batches ($m(t)$ and $M(t)$), the batched samples $\{X_{k}\}_{k\in M(t)}$, and the resampled batch mean $\mu_{\star}(\tau)$. Recall that $m(0)=0$, m(t)>m(t1)m(t)>m(t-1), $M(t)=\{m(t-1)+1,\cdots,m(t)\}$ for all $t\in\mathbb{N}$, and $|M(t)|\to+\infty$ (Assumption 2.3). Now, suppose it turns out $\tau=5$ based upon the past and latest information $\mathcal{F}_{m(5)}$. The output of our stopping rules is not the empirical batch mean $\mu(5)(=|M(5)|^{-1}\sum_{k\in M(5)}X_{k})$ along the top line, but the resampled empirical batch mean $\mu_{\star}(5)$ over the $5$th batch along the second line carrying on the past information $\mathcal{F}_{m(4)}$ (indicted by the thick arrow). The resampled sequence $\{X_{k}^{\star}\}_{k\in M(\tau)}$ is conditionally independent of the original counterpart $\{X_{k}\}_{k\in M(\tau)}$ on the stopping-time $\sigma$-field $\mathcal{F}_{m(\tau)}$.
Example 3.1.
Here, we demonstrate the procedure for constructing the resampled empirical batch mean (3.1) along the concrete problem settings presented in Example 2.2. As illustrated in Figure 1, suppose it turns out $\tau=5$ along the sequence $\{X_{k}\}_{k\in\mathbb{N}}$.
(a) If $\{X_{k}\}_{k\in\mathbb{N}}$ is a sequence of iid random variables, then generate the sequence of iid random variables $\{X^{\star}_{k}\}_{k\in M(\tau)}$, each under the same law as $\mathcal{L}(X_{1})$ and independently of the original sequence $\{X_{k}\}_{k\in\mathbb{N}}$. This is exactly what is done in the existing stopping rules for the standard iid Monte Carlo simulation (for instance, [1]).
(b) If $\{X_{k}\}_{k\in\mathbb{N}}$ is an ARCH(1) sequence with $X_{k}=(\beta+\alpha(X_{k-1})^{2})^{1/2}V_{k}$, then construct the sequence $\{X^{\star}_{k}\}_{k\in M(\tau)}$, as follows:
$X^{\star}_{k}=\begin{dcases}(\beta+\alpha(X_{k-1})^{2})^{1/2}V^{\star}_{k},&\text{if }k=m(\tau-1)+1,\\ (\beta+\alpha(X^{\star}_{k-1})^{2})^{1/2}V^{\star}_{k},&\text{if }k\in\{m(\tau-1)+2,\cdots,m(\tau)\},\end{dcases}$
[…]
(c) In the context of adaptive control variates and importance sampling, generate the sequence $\{U^{\star}_{k}\}_{k\in M(\tau)}$ of iid uniform random variables on $(0,1)$, independent of the original sequence $\{U_{k}\}_{k\in M(\tau)}$. Accordingly, in a similar manner to the construction of the augmented $\sigma$-field $\mathcal{F}^{\star}_{k}$, branch the natural filtration $(\mathcal{F}_{k})_{k\in\{0,1,\cdots,m(\tau-1)\}}$ on the basis of the resampled sequence $\{U^{\star}_{k}\}_{k\in M(\tau)}$, as well as construct $\{{\bm{\theta}}^{\star}_{k}\}_{k\in M(\tau)}$ in an adaptive manner to this branched filtration. Then, construct
$X^{\star}_{k}=\begin{dcases}R(U^{\star}_{k};{\bm{\theta}}_{k-1}),&\text{if }k=m(\tau-1)+1,\\ R(U^{\star}_{k};{\bm{\theta}}^{\star}_{k-1}),&\text{if }k\in\{m(\tau-1)+2,\cdots,m(\tau)\},\end{dcases}$
[…]
The motivation for introducing the resampled empirical batch mean $\mu_{\star}(\tau)$ is nothing but its unbiasedness while the one $\mu(\tau)$ based on the original sequence $\{X_{k}\}_{k\in\mathbb{N}}$ can be biased. To be exact, although the stopping batch $\tau$ is determined based on the original sequence $\{X_{k}\}_{k\in\mathbb{N}}$ (the top line of Figure 1), the output needs to be the resampled empirical batch mean $\mu_{\star}(\tau)$ at the stopping batch $\tau$ to ensure unbiasedness, as formalized below.
Theorem 3.2.
It holds $\mathbb{P}_{0}$-$a.s.$ that $\mathbb{E}_{m(\tau)}[\mu_{\star}(\tau)]=\mu$.
4 The stopping rules
The main idea of the present work is to replace the assessment of the error probability $\mathbb{P}_{0}(|\mu(t)-\mu|>\epsilon)$ with more implementable quantities for constructing a practical stopping rule. Along this line, such a quantity that we first work with is the (scaled) theoretical batch variance and standard deviation, respectively, by
$v^{2}_{0}(t)$ $:=|M(t)|{\rm Var}_{0}(\mu(t))=\frac{1}{|M(t)|}\sum_{k\in M(t)}{\rm Var}_{0}\left(X_{k}\right)$
$=\frac{1}{|M(t)|}\sum_{k\in M(t)}\mathbb{E}_{0}\left[X_{k}^{2}\right]-\mu^{2},$ (4.1)
and $v_{0}(t):=\left(v^{2}_{0}(t)\right)^{1/2}$ for $t\in\mathbb{N}$, where the first equality holds true by Theorem 2.4 (i) and the tower property of conditional expectation. Note that the quantities (4.1) are still not implementable in general, whereas they play an important role in our development. We first impose the following assumptions on its asymptotic behavior.
Assumption 4.1.
In addition to Assumptions 2.1 and 2.3, we impose $\inf_{t\in\mathbb{N}}v^{2}_{0}(t)>0$, $\sup_{t\in\mathbb{N}}v^{2}_{0}(t)<+\infty$ and $v^{2}_{0}(t+1)/|M(t+1)|\sim v^{2}_{0}(t)/|M(t)|$ as $t\to+\infty$, throughout.
Plainly speaking, the three assumptions here describe that over every single batch, its sample is neither completely degenerate nor exploding in variance, as well as the batch variance stays in a common order in the long run. It is worth noting that Theorem 2.4 (ii) holds true under the second assumption ($\sup_{t\in\mathbb{N}}v^{2}_{0}(t)<+\infty$) here, since the condition ($\limsup_{t\to+\infty}\sum_{k\in M(t)}(k-m(t-1))^{-2}{\rm Var}_{0}(X_{k})<+\infty$) in Theorem 2.4 (ii) is weaker. For instance, with ${\rm Var}_{0}(X_{k})\sim(k-m(t-1))^{q}$ for any $q\in(0,1)$, we have $v^{2}_{0}(t)\to+\infty$ while the condition $\limsup_{t\to+\infty}\sum_{k\in M(t)}(k-m(t-1))^{-2}{\rm Var}_{0}(X_{k})<+\infty$ holds true. Hence, the convergence $\lim_{t\to+\infty}\mu(t)=\mu$ remains true $\mathbb{P}_{0}$-$a.s.$, in what follows. The third assumption ($v^{2}_{0}(t+1)/|M(t+1)|\sim v^{2}_{0}(t)/|M(t)|$) does not look intuitive but is rather technical for asymptotic validation of the proposed stopping rules. Here, we present a standard form of the martingale central limit theorem [25] in our notation.
Theorem 4.2.
If $|M(t)|^{-1}\sum_{k\in M(t)}\mathbb{E}_{k-1}[|X_{k}-\mu|^{2}\mathbbm{1}(|X_{k}-\mu|>\epsilon\sqrt{|M(t)|})]\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}0$ as $t\to+\infty$ for all $\epsilon>0$, then it holds under $\mathbb{P}_{0}$ that
$\sqrt{|M(t)|}\frac{\mu(t)-\mu}{v_{0}(t)}\stackrel{{\scriptstyle\mathcal{L}}}{{\to}}\mathcal{N}(0,1),$ (4.2)
as $t\to+\infty$.
4.1 The stopping rule based on theoretical batch variances
In light of the weak convergence (4.2), we start with the following stopping rule and its approximation with “no” view towards implementation:
$\tau^{0}_{\epsilon,\delta}:=\mathop{\rm argmin}_{t\in\mathbb{N}}\left\{\mathbb{P}_{0}(|\mu(t)-\mu|>\epsilon)\leq\delta\right\},$ (4.3)
with precision $\epsilon\in(0,+\infty)$ and error probability $\delta\in(0,1)$ specified in subscript for asymptotic justification, unlike in Section 3. As is obvious from its expression, the stopping batch $\tau_{\epsilon,\delta}^{0}$ is $\mathcal{F}_{0}$-measurable, that is, deterministic in effect, as well as is far from implementable because the law of $|\mu(t)-\mu|$ is unknown. Still, at this initiation phase, we justify its “asymptotic” relevance in a similar manner to [12], as if the precision $\epsilon$ and the error probability $\delta$ could be adjusted after the sequence $\{X_{k}\}_{k\in\mathbb{N}}$ of infinite length had been generated.
Proposition 4.3.
(i) It holds that $\tau^{0}_{\epsilon,\delta}\to+\infty$ as $\epsilon\to 0+$ or $\delta\to 0+$.
For (ii) and (iii), let the conditions of Theorem 4.2 hold and fix $\delta\in(0,1)$.
(ii) It holds that $2(1-\Phi(\epsilon\sqrt{|M(\tau^{0}_{\epsilon,\delta})|}/v_{0}(\tau^{0}_{\epsilon,\delta})))\to\delta$, as $\epsilon\to 0+$.
(iii) If $v^{2}_{0}(t)/|M(t)|\sim ct^{-2q}$ as $t\to+\infty$ for some positive $c$ and $q$, then it holds that $\epsilon^{1/q}\tau^{0}_{\epsilon,\delta}\to(\sqrt{c}\Phi^{-1}(1-\delta/2))^{1/q}$ and $\epsilon^{-1}(\mu(\tau^{0}_{\epsilon,\delta})-\mu)\stackrel{{\scriptstyle\mathcal{L}}}{{\to}}\mathcal{N}(0,1/(\Phi^{-1}(1-\delta/2))^{2})$, as $\epsilon\to 0+$.
In short, it is confirmed in Part (i) that the simulation needs to go on forever if the criterion (4.3) is conducted rigidly ($\epsilon\to 0+$ or $\delta\to 0+$), just as it should be in principle. Moreover, let us emphasize the divergence $\tau^{0}_{\epsilon,\delta}\to+\infty$ here is purely deterministic, with nothing to do with probabilistic convergence modes, such as almost sure or in probability. Part (ii) justifies the replacement of the stopping rule $\tau_{\epsilon,\delta}^{0}$ by the following approximation in reference to the weak convergence (4.2), for sufficiently small precision $\epsilon$:
$\tau^{0}_{\epsilon,\delta}$ $=\mathop{\rm argmin}_{t\in\mathbb{N}}\left\{\mathbb{P}_{0}(|\mu(t)-\mu|>\epsilon)\leq\delta\right\}$
$\sim\mathop{\rm argmin}_{t\in\mathbb{N}}\left\{2\left(1-\Phi\left(\frac{\epsilon\sqrt{|M(t)|}}{v_{0}(t)}\right)\right)\leq\delta\right\},$ (4.4)
where the fraction inside the standard normal distribution function is implementable for all $t\in\mathbb{N}$ due to $\inf_{t\in\mathbb{N}}v^{2}_{0}(t)>0$ (Assumption 4.1). Indeed, the effectiveness of this approximation is the key building block for all the following developments in the upcoming sections. As for Part (iii), the first result indicates that the stopping batch increases in the order of $\epsilon^{-1/q}$ as the precision $\epsilon$ tends to vanish, depending on the exponent $q$ in the assumption $v^{2}_{0}(t)/|M(t)|\sim ct^{-2q}$. In the context of typical non-batched experiments with empirical mean $n^{-1}\sum_{k=1}^{n}X_{k}$ and convergent empirical variance, this assumption often holds with $q=1/2$, resulting in the asymptotic behavior of the stopping rule $\tau^{0}_{\epsilon,\delta}$ like $\epsilon^{-2}$. Next, the second result of Part (iii) describes the second-order structure of the output $\mu(\tau_{\epsilon,\delta}^{0})$ with the limiting variance $1/(\Phi^{-1}(1-\delta/2))^{2}$. That is, the experiment is expected to be more precise (that is, a smaller limiting variance) when the value of the error probability $\delta$ is very small, just as desired.
In turn, for implementation purposes, that is, in the non-asymptotic regime with both $\epsilon$ and $\delta$ fixed, the stopping rule and its approximation (4.4) need to be reformulated in a way to approximate the failure probability $\mathbb{P}_{0}(|\mu(t)-\mu|>\epsilon)$ by the conditional probability $\mathbb{P}_{m(\tau^{0}_{\epsilon,\delta})}(|\mu_{\star}(\tau^{0}_{\epsilon,\delta})-\mu|>\epsilon)$, in the sense of
$\mathbb{P}_{m(\tau^{0}_{\epsilon,\delta})}\left(|\mu_{\star}(\tau^{0}_{\epsilon,\delta})-\mu|>\epsilon\right)$
$\quad=\mathbb{P}_{m(\tau^{0}_{\epsilon,\delta})}\left(\sqrt{|M(\tau^{0}_{\epsilon,\delta})|}\frac{|\mu_{\star}(\tau^{0}_{\epsilon,\delta})-\mu|}{v(\tau^{0}_{\epsilon,\delta})}>\sqrt{|M(\tau^{0}_{\epsilon,\delta})|}\frac{\epsilon}{v(\tau^{0}_{\epsilon,\delta})}\right)$
$\quad\approx 2\left(1-\Phi(\epsilon\sqrt{|M(\tau^{0}_{\epsilon,\delta})|}/v(\tau^{0}_{\epsilon,\delta}))\right),$ (4.5)
provided that the value of the stopping batch $\tau_{\epsilon,\delta}^{0}$ is luckily far away. We stress that the last approximation in (4.5) is valid because the resampled batch $\{X_{k}^{\star}\}_{k\in M(\tau)}$ is conditionally independent of the stopping batch $\tau_{\epsilon,\delta}^{0}$ on the stopping-time $\sigma$-field $\mathcal{F}_{m(\tau_{\epsilon,\delta}^{0})}$.
4.2 The stopping rule based on implementable batch variances
As is now obvious from its expression, the crucial drawback of the approximate stopping rule (4.5) is the unavailability of the theoretical batch standard deviation $\{v_{0}(t)\}_{t\in\mathbb{N}}$. In what follows, we aim to replace the theoretical batch standard deviation in the approximate criterion (4.4) with a more implementable sequence $\{v(t)\}_{t\in\mathbb{N}}$ and make the resulting criterion more precise for a better application of the following martingale central limit theorem [27, Chapter VIII.1], which is a direct consequence of Theorem 4.2 owning to the continuous mapping theorem.
Theorem 4.4.
Let $\{v(t)\}_{t\in\mathbb{N}}$ be an $(\mathcal{F}_{m(t)})_{t\in\mathbb{N}}$-adapted sequence of random variables satisfying $\inf_{t\in\mathbb{N}}v^{2}(t)>0$, $\sup_{t\in\mathbb{N}}v^{2}(t)<+\infty$ and $v^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}1$, as $t\to+\infty$, and let the conditions of Theorem 4.2 hold. Then, it holds under $\mathbb{P}_{0}$ that
$\sqrt{|M(t)|}\frac{\mu(t)-\mu}{v(t)}\stackrel{{\scriptstyle\mathcal{L}}}{{\to}}\mathcal{N}(0,1),$ (4.6)
as $t\to+\infty$.
In the literature, a variety of sufficient conditions have been provided in different forms for the weak convergence (4.6) to hold true, which we do not pursue as those are not directly relevant in our context. For detail, we refer the reader to, for instance, [16, Theorem 2.5]. Let us stress that the condition $v^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}1$ implies that the sequence $\{v(t)\}_{t\in\mathbb{N}}$ does not vanish in the limit, due to Assumption 4.1. From a practical point of view, this is a reasonable property, since the sample $\{X_{k}\}_{k\in\mathbb{N}}$ itself (not the sample average) satisfying $v^{2}(t)\to 0$ would otherwise necessarily converge to the constant $\mu$ of interest, for which a stopping rule would no longer be of particular use.
With the sequence $\{v(t)\}_{t\in\mathbb{N}}$ (in place to the impractical one $\{v_{0}(t)\}_{t\in\mathbb{N}}$), the approximate stopping rule (4.4) can be advanced a step further to the following stopping rule:
$\tau_{\epsilon,\delta}:=\mathop{\rm argmin}_{t\in\mathbb{N}}\left\{2\left(1-\Phi\left(\frac{\epsilon\sqrt{|M(t)|}}{v(t)+a(t)}\right)\right)\leq\delta\right\},$ (4.7)
where $a(t)$ denotes a strictly positive function that decreases monotonically to zero as $t\to+\infty$, equipped for the purpose of asymptotic justification for the moment. Later in Section 5, we touch on the effect of the term $a(t)$ on the performance of the stopping rule (4.7).
Let us emphasize that the sequence $\{v(t)\}_{t\in\mathbb{N}}$ in (4.7) cannot be built on the resampled sequence $\{X_{k}^{\star}\}_{k\in M(\tau_{\epsilon,\delta})}$, because $\tau_{\epsilon,\delta}$ needs to be an $(\mathcal{F}_{m(t)})_{t\in\mathbb{N}_{0}}$-stopping time, generated by the original sequence $\{X_{k}\}_{k\in\mathbb{N}}$ alone. On the basis of the stopping time (4.7), we construct a stopping rule as follows.
Input: Precision $\epsilon$, error probability $\delta$, random seed $\omega$ and batch indices $\{m(t)\}_{t\in\mathbb{N}_{0}}$
Output: $\tau_{\epsilon,\delta}(\omega)$ and $\mu_{\star}(\tau_{\epsilon,\delta})(\omega)$
$t\leftarrow 1$;
while $2(1-\Phi(\epsilon\sqrt{|M(t)|}/(v(t)+a(t))))(\omega)>\delta$ do
 $t\leftarrow t+1$;
Set $\tau_{\epsilon,\delta}(\omega)\leftarrow t$ and compute $\mu_{\star}(\tau_{\epsilon,\delta})(\omega)$.
Algorithm 1 Stopping rule based on a more implementable batch variance
In a similar manner to [3], the “asymptotic” relevance of the criterion (4.7) can be justified in a way that the precision $\epsilon$ and the error probability $\delta$ are adjusted after the sequence $\{X_{k}\}_{k\in\mathbb{N}}$ of infinite length (which is certainly impossible to keep in hand) has been generated. We note that the almost sure convergence $v^{2}(t)/v^{2}_{0}(t)\to 1$ can be weakened to the weak convergence $\{v^{2}(\lfloor t/\epsilon\rfloor)/v^{2}_{0}(\lfloor t/\epsilon\rfloor):\,t\in(0,+\infty)\}\to 1$ in $\mathbb{D}((0,+\infty);\mathbb{R})$ as $\epsilon\to 0+$ for a version of Proposition 4.5 based on the functional weak law of large numbers, while it cannot be weakened further to $v^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}1$ since then the corresponding version of Part (ii) would not hold anymore (see, for instance, [13]).
Proposition 4.5.
Suppose $\lim_{t\to+\infty}v^{2}(t)/v^{2}_{0}(t)=1$, $\mathbb{P}_{0}$-$a.s.$, in addition to the conditions imposed in Theorem 4.4.
(i) It holds $\mathbb{P}_{0}$-$a.s.$ that $\tau_{\epsilon,\delta}\to+\infty$ as $\epsilon\to 0+$ or $\delta\to 0+$.
For (ii) and (iii), fix $\delta\in(0,1)$.
(ii) It holds $\mathbb{P}_{0}$-$a.s.$ that $2(1-\Phi(\epsilon\sqrt{|M(\tau_{\epsilon,\delta})|}/(v(\tau_{\epsilon,\delta})+a(\tau_{\epsilon,\delta})))\to\delta$, as $\epsilon\to 0+$.
(iii) If $v^{2}_{0}(t)/|M(t)|\sim ct^{-2q}$ as $t\to+\infty$ for some positive $c$ and $q$, then it holds under $\mathbb{P}_{0}$ that $\epsilon^{1/q}\tau_{\epsilon,\delta}\to(\sqrt{c}\Phi^{-1}(1-\delta/2))^{1/q}$ $a.s.$, and $\epsilon^{-1}(\mu_{\star}(\tau_{\epsilon,\delta})-\mu)\stackrel{{\scriptstyle\mathcal{L}}}{{\to}}\mathcal{N}(0,1/(\Phi^{-1}(1-\delta/2))^{2})$, as $\epsilon\to 0+$.
In relation to Proposition 4.5 (ii), we stress that $\mathbb{P}_{0}(|\mu(\tau_{\epsilon,\delta})-\mu|>\epsilon)\not\to\delta$ in general, as the random variable $\mu(\tau_{\epsilon,\delta})$ is not identical in law to the empirical batch mean $\mu(t)$ at a usual deterministic batch $t$ in the presence of the stopping time $\tau_{\epsilon,\delta}$. Conversely, the conditional standard deviation $s(\tau_{\epsilon,\delta})$ at the stopping batch $\tau_{\epsilon,\delta}$ cannot be replaced by its resampled version, that is, $v_{\star}^{2}(\tau_{\epsilon,\delta})=|M(\tau_{\epsilon,\delta})|^{-1}\sum_{k\in M(\tau_{\epsilon,\delta})}{\rm Var}_{k-1}^{\star}(X^{\star}_{k})$, since the resampled batch $\{X^{\star}_{k}\}_{k\in M(\tau)}$ is conditionally independent of the stopping batch $\tau_{\epsilon,\delta}$ on the stopping-time $\sigma$-field $\mathcal{F}_{m(\tau_{\epsilon,\delta})}$.
With Proposition 4.5 being presented, of practical interest is the non-asymptotic regime [1, 17] in which one generates a segment $\{X_{k}\}_{k\in M(t)}$ of the sequence, one batch at a time forward until, and only until, the criterion (4.7) is satisfied, with precision $\epsilon$ and error probability $\delta$ fixed.
4.2.1 Conditional batch variances
A candidate for the sequence $\{v(t)\}_{t\in\mathbb{N}}$ is based on the conditional batch variance and standard deviation, defined as
$v_{1}^{2}(t):=\frac{1}{|M(t)|}\sum_{k\in M(t)}{\rm Var}_{k-1}\left(X_{k}\right)=\frac{1}{|M(t)|}\sum_{k\in M(t)}\mathbb{E}_{k-1}\left[X_{k}^{2}\right]-\mu^{2},$ (4.8)
and $v_{1}(t):=\left(v_{1}^{2}(t)\right)^{1/2}$ for $t\in\mathbb{N}$, where we have applied the base assumption (2.1) for the evaluation of the conditional variance. The conditional batch variance here is expected to behave in a somewhat similar manner to the theoretical one in the sense that it is unbiased for the theoretical batch variance (4.1), that is, $\mathbb{E}_{0}[v_{1}^{2}(t)]=v^{2}_{0}(t)$ for all $t\in\mathbb{N}$. Based on the availability of the sample $\{X_{k}\}_{k\in M(t)}$ alone, the conditional batch variance and standard deviation (4.8) are, in principle, more implementable than the theoretical ones (4.1), whereas the reality is that neither is generally implementable, with a few exceptions (such as the one examined in Section 5.1, where the conditional one (4.8) happens to be available in an implementable form).
For later use, we amend the criterion (4.7) with $v(t)$ replaced by the conditional counterpart $v_{1}(t)$, as follows:
$\tau^{a}_{\epsilon,\delta}:=\mathop{\rm argmin}_{t\in\mathbb{N}}\left\{2\left(1-\Phi\left(\frac{\epsilon\sqrt{|M(t)|}}{v_{1}(t)+a(t)}\right)\right)\leq\delta\right\},$ (4.9)
as a stopping batch index.
4.2.2 Empirical batch variances
A more realistic candidate for the sequence $\{v(t)\}_{t\in\mathbb{N}}$ can be constructed based on the empirical batch variance and standard deviation, respectively, defined as
$v_{2}^{2}(t):=\frac{1}{|M(t)|-1}\sum_{k\in M(t)}\left(X_{k}-\mu(t)\right)^{2},\quad v_{2}(t):=\left(v_{2}^{2}(t)\right)^{1/2},\quad t\in\mathbb{N},$ (4.10)
where $\mu(t)$ is the (implementable) empirical batch mean (2.2), rather than the (unknown) constant $\mu$ in the base assumption (2.1). The relevance is that as opposed to the conditional batch variance and standard deviation (4.8), the empirical batch variance $v_{2}^{2}(t)$ and standard deviation $v_{2}(t)$ here are implementable as soon as the sample $\{X_{k}\}_{k\in M(t)}$ is given, yet with unbiasedness $\mathbb{E}_{0}[v_{2}^{2}(t)]=v^{2}_{0}(t)$ holding true. An empirical version of the martingale central limit theorem can be constructed on the basis of the empirical standard deviation (4.10), as follows.
Theorem 4.6.
In addition to the conditions imposed in Theorem 4.4, let the following two conditions hold: $\limsup_{k\to+\infty}\mathbb{E}_{0}[|X_{k}|^{4}]<+\infty$ and $\lim_{t\to+\infty}|M(t)|^{-1}\sum_{k\in M(t)}(X_{k}^{2}-\mathbb{E}_{k-1}[X_{k}^{2}])=0$, $\mathbb{P}_{0}\mbox{-}a.s.$ It then holds that $v_{2}^{2}(t)/v_{1}^{2}(t)\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}1$.
For later use, we amend the criterion (4.7) with $v(t)$ replaced by the empirical counterpart $v_{2}(t)$, as follows:
$\tau^{b}_{\epsilon,\delta}:=\mathop{\rm argmin}_{t\in\mathbb{N}}\left\{2\left(1-\Phi\left(\frac{\epsilon\sqrt{|M(t)|}}{v_{2}(t)+a(t)}\right)\right)\leq\delta\right\},$ (4.11)
as a stopping batch index.
5 Numerical illustrations
In this section, we present numerical examples to demonstrate the procedure and illustrate the effectiveness and the strong potential of the proposed stopping rules in making a satisfactory stop of the sequential experiment. To this end, we first summarize the developed stopping rules.
For precision $\epsilon\in(0,+\infty)$ and error probability $\delta\in(0,1)$, one stops the experiment at a batch $\tau_{\epsilon,\delta}$ in which the probability of the undesirable event $\{|\mu(t)-\mu|>\epsilon\}$ is satisfactorily small for the first time, that is, $\tau_{\epsilon,\delta}=\min_{t\in\mathbb{N}}\{\mathbb{P}_{m(t-1)}(|\mu(t)-\mu|>\epsilon)\leq\delta\}$. Depending on how to approximate the error probability $\mathbb{P}_{m(t-1)}(|\mu(t)-\mu|>\epsilon)$ in the criterion, we have developed the following two stopping rules:
[…]
where $v_{1}(t)$ and $v_{2}(t)$ are the conditional and empirical batch standard deviations, defined respectively in (4.8) and (4.10). To assess the proposed stopping rules, we take the following steps.
Fix the precision $\epsilon\in(0,+\infty)$ and the error probability $\delta\in(0,1)$.
Fix the batches $\{m(t)\}_{t\in\mathbb{N}}$, the number of iid experiments $n$, and then $n$ distinct random seeds $(\omega_{1},\cdots,\omega_{n})(=:\omega)$.
For every random seed $\omega_{k}$, obtain the stopping batch index $\tau_{\epsilon,\delta}(\omega_{k})$ in accordance with either (4.9) or (4.11), and compute the indicator $\mathbbm{1}(|\mu_{\star}(\tau_{\epsilon,\delta})(\omega_{k})-\mu|\leq\epsilon)$ over the $\tau_{\epsilon,\delta}(\omega_{k})$th batch in accordance with (3.1).
Compute an empirical error probability $p_{\epsilon,\delta}(\omega)$, the so-called reliability $R_{\epsilon,\delta}(\omega)$, and the complexity ${\rm CM}_{\epsilon,\delta}(\omega)$, respectively, as follows:
$p_{\epsilon,\delta}(\omega):=\frac{1}{n}\sum_{k=1}^{n}\mathbbm{1}\left(|\mu_{\star}(\tau_{\epsilon,\delta})(\omega_{k})-\mu|\leq\epsilon\right),\quad R_{\epsilon,\delta}(\omega):=\frac{p_{\epsilon,\delta}(\omega)}{1-\delta},$
$\text{CM}_{\epsilon,\delta}(\omega):=\frac{1}{n}\sum_{k=1}^{n}\left[(\tau_{\epsilon,\delta}(\omega_{k}))^{\ell}-({\tau_{\epsilon,\delta}(\omega_{k})}-1)^{\ell}\right].$ (5.2)
In accordance with [1], we consider a stopping rule with its reliability $R_{\epsilon,\delta}(\omega)$ strictly less than $1$ unreliable. In addition, the effectiveness of a stopping rule improves as its complexity $\text{CM}_{\epsilon,\delta}(\omega)$ decreases at the specified precision and error probability. We note that the parameter $\ell$ in the complexity $\text{CM}_{\epsilon,\delta}(\omega)$ controls how severely we penalize large stopping times and is suppressed from its notation because we fix $\ell=5$ throughout. To ensure sufficient realizations, we set $n=5\times 10^{3}$ above.
With suitable constants $0<a_{1}<a_{2}<+\infty$ and $0<b_{1}<b_{2}<1$, we split the rectangle $a_{1},a_{2}]\times[b_{1},b_{2}]$ into an logarithmically spaced grid of $10\times 10$ points, denoted by $\{(\epsilon_{k},\delta_{k})\}_{k\in\{1,\cdots,10\times 10\}}$. On each of these $10\times 10$ combinations of precision and error probability, we compute the reliability $R_{\epsilon_{k},\delta_{k}}(\omega)$ and the complexity ${\rm CM}_{\epsilon_{k},\delta_{k}}(\omega)$, with the random seeds $\omega$ fixed throughout.
5.1 ARCH(1)
We begin our numerical illustrations with an ARCH(1) model, which provides an ideal test case for our stopping rules due to its non-trivial dependence structure and heavy-tailed behavior, as well as the straightforward implementation of the theoretical and conditional batch variances. This setting poses natural challenges for sequential sampling methods, as the underlying dynamics combine temporal dependence through the ARCH structure with potentially extreme observations from the $t$ distribution.
Now, consider an ARCH(1) sequence $\{X_{k}\}_{k\in\mathbb{N}}$, which is generated recursively by $X_{k}=(\beta+\alpha X_{k-1}^{2})^{1/2}V_{k}$, with $X_{0}=0$, $\alpha\in(0,1)$ and $\beta\in(0,+\infty)$. Here, $\{V_{k}\}_{k\in\mathbb{N}}$ denotes a sequence of iid random variables following a scaled $t$-distribution with $n$ degrees of freedom ($n\in\{5,6,\cdots\}$), normalized to have unit variance, with its probability density function given as
$\frac{d}{dx}\mathbb{P}(V_{1}\leq x)=\frac{\Gamma((n+1)/2)}{\sqrt{(n-2)\pi}\Gamma(n/2)}\left(1+\frac{x^{2}}{n-2}\right)^{-(n+1)/2},\quad x\in\mathbb{R}.$
The theoretical and conditional batch variances for this model are given by
$v^{2}_{0}(t)=\frac{\beta}{1-\alpha}\left(1-\frac{\alpha^{m(t-1)+1}-\alpha^{m(t)}}{2}\right),\quad v_{1}^{2}(t)=\beta+\frac{\alpha}{|M(t)|}\sum_{k\in M(t)}X_{k-1}^{2}.$
The filtration $(\mathcal{F}_{k})_{k\in\mathbb{N}_{0}}$ is constructed as $\mathcal{F}_{k}:=\sigma(\mathcal{N})\vee\sigma(V_{1},\cdots,V_{k})$ for all $k\in\mathbb{N}_{0}$, while the branched filtration $(\mathcal{F}_{k-1}^{\star})_{k\in M(\tau)}$ is constructed in the way of $\mathcal{F}_{k-1}^{\star}=\mathcal{F}_{m(\tau)}\vee\sigma(V_{m(\tau-1)+1}^{\star},\cdots,V_{k-1}^{\star})$ for all $k\in M(\tau)$. We note that $\sigma$ here denotes the $\sigma$-field, rather than the empirical batch variance (4.10).
In what follows, we fix $\alpha=0.03$, $\beta=0.3$, and $n=6$ degrees of freedom. The batch structure is set as $M(t)=\{(t-1)^{5}+1,\cdots,t^{5}\}$, ensuring sufficient sample growth while maintaining theoretical requirements.
Figure 2: Typical realization of theoretical batch variance $v_{0}^{2}(t)$ (solid blue), conditional batch variance $v_{1}^{2}(t)$ (dashed red), and empirical batch variance $v_{2}^{2}(t)$ (dotted green) of the ARCH(1) process under consideration.
As we observe from Figure 2, after some initial fluctuations due to smaller batch sizes, the theoretical, conditional, and empirical variances exhibit similar behavior, with all those measures tracking each other closely. This validates the use of the empirical variance $v_{2}^{2}(t)$ in the stopping rule, as it succeeds in approximating the theoretical variance, not to mention the conditional variance, in the long run.
We implement $\tau^{b}_{\epsilon,\delta}$ based on the empirical batch variance $v_{2}^{2}(t)$ as of (5.1), as it can be readily available unlike the theoretical and conditional variances. Here, we set $a(t)=1/t$, which helps to compensate the severe underestimation of the empirical variance $v_{2}^{2}(t)$ during early batches, such as $v_{2}^{2}(1)\approx 0$ observed in Figure 2. The choice of $a(t)=1/t$ satisfies its theoretical requirement of monotonic vanishing and reduces the risk of premature stopping.
We assess its performance by examining reliability and computational complexity ((iv)) across various precision ($\epsilon$) and error probability ($\delta$) values. For each $(\epsilon_{k},\delta_{k})$ in the specified grid, we perform independent simulations to compute $\tau^{b}_{\epsilon_{k},\delta_{k}}(\omega_{j})$, generate new batches for the resampled mean $\mu_{\star}(\tau_{\epsilon_{k},\delta_{k}})(\omega_{j})$, and evaluate success through $\mathbbm{1}(|\mu_{\star}(\tau_{\epsilon_{k},\delta_{k}})(\omega_{j})-\mu|\leq\epsilon_{k})$.
Figure 3: The left plot represents the reliability $R_{\epsilon,\delta}$ and the right plot shows the complexity ${\rm CM}_{\epsilon,\delta}$ on a log10 scale.
Mean Min Max
Reliability ($R_{\epsilon,\delta}$) 1.0225 1.0010 1.1111
Complexity ($\rm CM_{\epsilon,\delta}$) 1,601,368.29 456.92 16,417,500.00
Table 1: Performance metrics
In Figure 3 and Table 1 above, we present our comprehensive numerical findings on the performance of the proposed stopping rule. The reliability plot (the left plot of Figure 3) demonstrates that our stopping rules consistently achieve the reliability $R_{\epsilon,\delta}$ staying strictly greater than $1$ across all parameter combinations. As confirmed in Table 1, reliability values range from 1.001 to 1.111, with a mean of 1.0225, satisfying the reliability criterion set up in [1]. The reliability increases systematically with larger values of $\epsilon$ and $\delta$, reflecting more conservative behavior in less demanding scenarios while maintaining reliability even in the most challenging case with $\epsilon=\delta=10^{-3}$. The complexity heat map (the right plot of Figure 3) reveals a pattern of computational cost across the parameter space. The logarithmic scale clearly shows how the complexity $\rm CM_{\epsilon,\delta}$ increases as both $\epsilon$ and $\delta$ decrease. The computational burden varies substantially, as detailed in Table 1, ranging from $\rm CM_{\epsilon,\delta}\approx 4.57\times 10^{2}$ for larger values of $\epsilon$ and $\delta$ to $\rm CM_{\epsilon,\delta}\approx 1.64\times 10^{7}$ for $\epsilon=\delta=10^{-3}$, with a mean of approximately 1.6 million samples. This scaling aligns with our theoretical predictions and provides practical guidance for computational requirements. Smooth transitions in both reliability and complexity across the parameter space indicate the stability and robustness of the proposed stopping rule. The consistent achievement of the requirement $R_{\epsilon,\delta}>1$, coupled with theoretically predicted complexity scaling, provides strong empirical validation of our methodology, particularly notable given the challenging features of temporal dependence and heavy-tailed distributions in the ARCH(1) setting.
5.2 Adaptive control variates
In this section, we examine the problem setting described in Examples 2.2 (c) and 3.1 (c). To recall it, suppose, for $\Psi:(0,1)^{d}\to\mathbb{R}$, we wish to evaluate the integral $\mu=\int_{(0,1)^{d}}\Psi({\bf u})d{\bf u}$, where the random element here is modeled via the uniform law on the unit hypercube $(0,1)^{d}$, which can be considered general enough thanks to the inverse transform random sampling method. Here, without affecting the value of the integral, one can introduce a parameter ${\bm{\theta}}\in\mathbb{R}^{d}$ in a linear manner, as follows
$\mu=\int_{(0,1)^{d}}\Psi({\bf u})d{\bf u}=\int_{(0,1)^{d}}\left(\Psi({\bf u})+\left\langle{\bm{\theta}},{\bf u}-\mathbbm{1}_{d}/2\right\rangle\right)d{\bf u},$ (5.3)
under the quartic integrability condition $\int_{(0,1)^{d}}|\Psi({\bf u})|^{4}d{\bf u}<+\infty$ for technical reasons (Section A.2). Note that the control variate $U-\mathbbm{1}_{d}/2$ is centered with variance-covariance matrix $\int_{(0,1)^{d}}({\bf u}-\mathbbm{1}_{d}/2)^{\otimes 2}d{\bf u}=\mathbb{I}_{d}/12$, that is, the diagonal matrix of order $d$ with all diagonal elements of $1/12$. Although one may introduce nonlinear control variates, we do not go in that direction in the present paper. Now, define $V({\bm{\theta}}):=\int_{(0,1)^{d}}(\Psi({\bf u})+\langle{\bm{\theta}},{\bf u}-\mathbbm{1}_{d}/2\rangle)^{2}d{\bf u}$, which represents the second moment of the rightmost expression in (5.3), parameterized by $\bm{\theta}$}. The estimator variance $V({\bm{\theta}})-\mu^{2}$ is, due to its quadratic structure in the parameter $\bm{\theta}$}, minimized uniquely at
${\bm{\theta}}^{*}:=\mathop{\rm argmin}_{{\bm{\theta}}\in\mathbb{R}^{d}}V({\bm{\theta}})=-12\int_{(0,1)^{d}}\Psi({\bf u})({\bf u}-\mathbbm{1}_{d}/2)d{\bf u},$ (5.4)
with the minimized second moment
$\min_{{\bm{\theta}}\in\mathbb{R}^{d}}V({\bm{\theta}})=V({\bm{\theta}}^{*})=V(0)-\frac{1}{12}\|{\bm{\theta}}^{*}\|^{2},$ (5.5)
where the first term $V(0)$ represents the crude second moment. Hence, the minimized second moment (5.5) is strictly smaller than the crude second moment $V(0)$ unless ${\bm{\theta}}^{*}=0_{d}$, that is, the integrand $\Psi({\bf u})$ is linearly uncorrelated with the underlying random element $\bf u$}, which is rather absurd from a practical point of view. It is known that Monte Carlo averaging, albeit no longer iid due to the evolving adaptive sequence $\{{\bm{\theta}}_{k}\}_{k\in\mathbb{N}_{0}}$, can proceed concurrently along with the parameter search, as follows:
$\frac{1}{n}\sum_{k=1}^{n}\left(\Psi(U_{k})+\left\langle{\bm{\theta}}_{k-1},U_{k}-\mathbbm{1}_{d}/2\right\rangle\right)\to\mu,\quad{\bm{\theta}}_{n}:=-12\frac{1}{n}\sum_{k=1}^{n}\Psi(U_{k})(U_{k}-\mathbbm{1}_{d}/2)(\to{\bm{\theta}}^{*}),$ (5.6)
with ${\bm{\theta}}_{0}:=0_{d}$, where $\{U_{k}\}_{k\in\mathbb{N}}$ is a sequence of iid uniform random vectors on $(0,1)^{d}$. Note that both convergences in (5.6) hold true $\mathbb{P}_{0}$-$a.s.$ as $n\to+\infty$.
In the present context, instead of updating the parameter $\bm{\theta}$} at every iteration as in (5.6), computing budget may be saved by updating it only once at the end of each batch (right before moving on to the next batch), as follows:
${\bm{\theta}}(t):=-12\frac{1}{|M(t)|}\sum_{k\in M(t)}\Psi(U_{k})(U_{k}-\mathbbm{1}_{d}/2),\quad t\in\mathbb{N},$ (5.7)
with ${\bm{\theta}}(0)=0_{d}$. By representing ${\bm{\theta}}(t)$ as a function of the batch index, we stress that it is updated only once for each batch. It clearly remains true that ${\bm{\theta}}(t)\to{\bm{\theta}}^{*}$, $\mathbb{P}_{0}$-$a.s.$ as $t\to+\infty$, just as in (5.6), because this is still nothing but standard iid Monte Carlo methods for the rightmost expression of (5.4). Here, the filtration $(\mathcal{F}_{k})_{k\in\mathbb{N}_{0}}$ can be set as $\mathcal{F}_{k}=\sigma(\mathcal{N})\vee\sigma(U_{1},\cdots,U_{k})$ for $k\in\mathbb{N}$, as well as the branched one $(\mathcal{F}_{k-1}^{\star})_{k\in M(\tau)}$ by $\mathcal{F}_{k-1}^{\star}=\mathcal{F}_{m(\tau_{\epsilon,\delta})}\vee\sigma(U_{m(\tau_{\epsilon,\delta}-1)+1}^{\star},\cdots,U_{k-1}^{\star})$ for all $k\in M(\tau_{\epsilon,\delta})$ on the iid resampled batch $\{U_{k}^{\star}\}_{k\in M(\tau_{\epsilon,\delta})}$, where $\tau_{\epsilon,\delta}$ is a suitable stopping batch as in (5.1). Then, for every $t\in\mathbb{N}$ and $k\in M(t)$, the base formulation (Assumption 2.1) holds by setting $X_{k}=\Psi(U_{k})+\left\langle{\bm{\theta}}(t-1),U_{k}-\mathbbm{1}_{d}/2\right\rangle$, as
$\mathbb{E}_{k-1}\left[X_{k}\right]$ $=\mathbb{E}_{k-1}\left[\Psi(U_{k})+\left\langle{\bm{\theta}}(t-1),U_{k}-\mathbbm{1}_{d}/2\right\rangle\right]$
$=\mathbb{E}_{k-1}\left[\Psi(U_{k})\right]+\left\langle{\bm{\theta}}(t-1),\mathbb{E}_{k-1}\left[U_{k}-\mathbbm{1}_{d}/2\right]\right\rangle=\mu,$
since ${\bm{\theta}}(t-1)$ is $\mathcal{F}_{k-1}$-measurable and the uniform random vector $U_{k}$ is independent of the $\sigma$-field $\mathcal{F}_{k-1}$ for all $k\in M(t)$. Hence, we aim to construct stopping rules for the algorithm:
$\mu(t)$ $:=\frac{1}{|M(t)|}\sum_{k\in M(t)}\left(\Psi(U_{k})+\left\langle{\bm{\theta}}(t-1),U_{k}-\mathbbm{1}_{d}/2\right\rangle\right),$ (5.8)
$v_{2}^{2}(t)$ $:=\frac{1}{|M(t)|}\sum_{k\in M(t)}\left(\Psi(U_{k})+\left\langle{\bm{\theta}}(t-1),U_{k}-\mathbbm{1}_{d}/2\right\rangle\right)^{2}-\mu^{2}(t),$
along with (5.7), each of which only requires elementary operations.
For illustrative purposes, we examine a one-dimensional integration problem with integrand $\Psi(u)=u^{2}/2$ on $(0,1)$, which has theoretical mean $\mu=1/6$. Our analysis shows the crude Monte Carlo estimator has variance $V(0)=2.22\times 10^{-2}$, while the theoretically minimized variance with optimal control variate is $V(\theta^{*})=1.39\times 10^{-3}$ at $\theta^{*}=-1/2$.
Figure 4: We plot empirical means $\mu(n)$, showing convergence behavior of both crude Monte Carlo (blue) and control variates (green) estimators towards the theoretical mean $\mu=1/6$ (dotted red) in the leftmost, an evolution of the adaptive control variates parameter $\theta(n)$ with its limiting value $\theta^{*}=-1/2$ (dotted red) in the middle, and empirical variances $v^{2}_{2}(n)$, with (green) and without (blue) adaptive control variates, with their respective limiting values $V(0)=2.22\times 10^{-2}$ and $V(\theta^{*})=1.39\times 10^{-3}$ in the rightmost.
In Figure 4 above, we present numerical results based upon the cumulative sample size $n$ to offer a more continuous view of the convergence behavior, as opposed to the batch index $t$ used in our theoretical development. The plots provide several important insights, as follows. The control variate estimator (green) demonstrates notably faster convergence to the theoretical mean $\mu=1/6$ (dotted red), compared to the crude Monte Carlo estimator (blue), with visibly reduced fluctuations from early stages. The control variate parameter $\theta(n)$ shows initial volatility but steadily converges to the theoretical optimum $\theta^{*}=-0.5$ (dotted red). This convergence corresponds directly to the improved performance in mean estimation and variance reduction. The empirical variance plot clearly demonstrates the effectiveness of adaptive control variates, with the control variate estimator (green) achieving and maintaining significantly lower variance compared to the crude Monte Carlo counterpart (blue).
In Figure 5 and Table 2 below, we present comprehensive performance metrics across different precision and error probability parameters. The results demonstrate that our stopping rules maintain robust performance even in the presence of adaptive parameter estimation ($\theta(t)$) and the inherent correlation structure between successive batches. We again set $a(t)=1/t$ as the inflation factor in the stopping rule (4.11). The reliability consistently exceeds 1 (ranging from 1.0010 to 1.1111) despite the additional complexity introduced by the concurrent parameter optimization and variance reduction processes. The complexity metrics reveal relatively efficient computation compared to the ARCH(1) case, with a mean complexity of approximately 63,881 samples - about 25 times lower than the ARCH(1) scenario. This improved efficiency can be attributed to the variance reduction effect of the adaptive control variates, which progressively optimizes $\theta(t)$ to minimize the estimated variance. The complexity scaling remains predictable across the parameter space, ranging from about 252 samples for larger $\varepsilon$ and $\delta$ values to around 542,264 samples in the most demanding case ($\varepsilon=\delta=10^{-3}$), demonstrating that the stopping rule effectively balances computational cost with precision requirements even as the control variate parameters are being adaptively optimized.
Figure 5: The left plot represents the reliability $R_{\epsilon,\delta}$ and the right plot shows the complexity ${\rm CM}_{\epsilon,\delta}$ on a log10 scale.
Mean Min Max
Reliability ($R_{\epsilon,\delta}$) 1.0265 1.0010 1.1111
Complexity ($\rm CM_{\epsilon,\delta}$) 63880.9398 252.0000 542264.0200
Table 2: Performance metrics
Remark 5.1.
It is worth noting that the proposed stopping rules for adaptive control variates have an interesting connection to averaged stochastic gradient descent (ASGD) [2]. While conventional SGD implementations often rely on validation performance for early stopping [28], monitoring averaged iterates, which helps reduce the effect of the noise term, has proven beneficial in practice. This suggests potential value in developing statistically driven stopping criteria that leverage averaging, as in our framework, particularly for validation-free scenarios [5]. ∎
6 Concluding remarks
We have developed an easy-to-implement sequential stopping rule for Monte Carlo methods with a primal view towards the mean estimation of a non-iid sequence of martingale difference type in the non-asymptotic regime. In contrast to standard iid Monte Carlo methods, where the classical central limit theorem suffices, the present aim relies largely on the martingale central limit theorem, which is, however, impractical in its original for most examples. We have addressed this difficulty by carefully imposing a range of step-by-step technical conditions so that the corresponding stopping rules remain feasible. We have provided numerical results, illustrating the applicability and effectiveness of the proposed stopping rules.
We conclude this work by highlighting two promising directions for future research. The first involves improving the quality of the proposed stopping rules. Unlike stopping rules based on the classical central limit theorem (such as [1, 17]), a key challenge in enhancing the quality of stopping rules here arises from the lack of accurate finite-time approximations for the limiting law in the context of the martingale central limit theorem. Specifically, obtaining Berry-Esseen-type bounds based on higher-order moments, and moreover presenting them in an implementable form, remains out of reach. Although there is extensive literature on martingale central limit theorems and their rates of convergence - typically expressed as uniform bounds (such as [4, 7, 24]) and, in some cases, as non-uniform bounds (such as [14, 15, 19]) - these results generally involve unspecifiable constants, limiting their practical applicability to the development of stopping rules, with an exception [6], which, to the best of our knowledge, provides the only implementable result of the uniform bound of Berry-Esseen type.
The second direction focuses on applying the proposed stopping rules across various fields. In this study, we did not perform a comprehensive numerical analysis across a wide range of practical problem settings. Given the advent of machine learning methods as briefly described in the introductory section and Remark 5.1, it would be especially beneficial to adapt and apply the proposed stopping rules to these relevant contexts in a tailored manner. Those would be valuable future research directions, deserving of their own separate investigation.
7 Proofs
To maintain the flow of the paper, we collect proofs here. We try to keep the paper fairly self-contained, whereas we skip nonessential details of a somewhat routine nature in some instances to avoid overloading the Appendix. In line with Figure 1, recall that $\mathbb{E}^{\star}_{k-1}$ is the expectation taken under the probability measure $\mathbb{P}^{\star}_{k-1}$ on the $\sigma$-field $\mathcal{F}^{\star}_{k-1}$ that makes the branched sequence $(X_{1},\cdots,X_{m(\tau-1)},X^{\star}_{m(\tau-1)+1},\cdots,X^{\star}_{k-1})$ measurable for all $k\in M(\tau)$ and leaves the original stopping batch $(X_{m(\tau-1)+1},\cdots,X_{m(\tau)})$ out of its measurability.
Proof of Theorems 2.4 and 3.2.
Theorem 2.4 is elementary by the tower rule and the standing assumption $\mathbb{E}_{k-1}[X_{k}]=\mu$ (Assumption 2.1) for the unbiasedness, as well as due to, for instance, [8, VII.9] for the strong law of large numbers. For Theorem 3.2, the first result can be derived as
$\mathbb{E}_{m(\tau)}[\mu_{\star}(\tau)]$ $=\frac{1}{|M(\tau)|}\sum_{k\in M(\tau)}\mathbb{E}_{m(\tau)}\left[X_{k}^{\star}\right]=\frac{1}{|M(\tau)|}\sum_{k\in M(\tau)}\mathbb{E}_{m(\tau-1)}\left[X_{k}^{\star}\right]$
$=\frac{1}{|M(\tau)|}\sum_{k\in M(\tau)}\mathbb{E}_{m(\tau-1)}\left[\mathbb{E}_{k-1}^{\star}\left[X_{k}^{\star}\right]\right]=\frac{1}{|M(\tau)|}\sum_{k\in M(\tau)}\mathbb{E}_{m(\tau-1)}[\mu],$ (7.1)
where we have applied the $\mathcal{F}_{m(\tau)}$-measurability of the stopping batch $\tau$ for the first equality and the tower rule with $\mathbb{E}^{\star}_{k-1}[X_{k}^{\star}]=\mathbb{E}_{k-1}[X_{k}]=\mu$ (Assumption 2.1) for the third and fourth equalities, due to $\mathcal{F}_{m(\tau-1)}\subseteq\mathcal{F}^{\star}_{k-1}$ for all $k\in M(\tau)=\{m(\tau-1)+1,\cdots,m(\tau)\}$. The second result is then straightforward due to $\mathcal{F}_{0}\subseteq\mathcal{F}_{m(\tau)}$. ∎
Proof of Theorem 4.4.
A standard martingale central limit theorem (for instance, [16, Theorem 2.5 (a)] with $\xi_{t,k}=(X_{k}-\mu)/\sqrt{|M(t)|v^{2}_{0}(t)}$) reads in the present framework as follows. If $v_{1}^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}1$ and $(|M(t)|v^{2}_{0}(t))^{-1}\sum_{k\in M(t)}\mathbb{E}_{k-1}[|X_{k}-\mu|^{2}\mathbbm{1}(|X_{k}-\mu|>\epsilon\sqrt{|M(t)|}v_{0}(t))]\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}0$ for every $\epsilon>0$, then it holds that $\mathbb{P}_{0}(\sqrt{|M(t)|}(\mu(t)-\mu)/v_{0}(t)\leq x)\to\Phi(x)$ for all $x\in\mathbb{R}$, as $t\to+\infty$. This Lindeberg condition can be simplified as the one in Theorem 4.2, thanks to $\inf_{t\in\mathbb{N}}v^{2}_{0}(t)>0$ and $\sup_{t\in\mathbb{N}}v^{2}_{0}(t)<+\infty$ (Assumption 4.1). The desired weak convergence (4.6) holds true by the Slutsky theorem due to the condition $v_{1}^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}1$. ∎
Next, we review a standard result on the strong law of large numbers for a sequence of asymptotically pairwise uncorrelated random variables. This result has already been employed in Section 5.1, as well as will be used in the proof of Theorem 4.6.
Lemma 7.1.
Let $\{Y_{k}\}_{k\in\mathbb{N}}$ be a sequence of random variables, satisfying the following conditions: $\lim_{k\to+\infty}\mathbb{E}_{0}[Y_{k}](=:c)$ exists, $|\mathbb{E}_{0}[(Y_{k}-\mathbb{E}_{0}[Y_{k}])(Y_{m}-\mathbb{E}_{0}[Y_{m}])]|=\mathcal{O}(m^{-2})$ for all $k\in\mathbb{N}$, and $\sup_{k\in\mathbb{N}}{\rm Var}_{0}(Y_{k})<+\infty$. Then, it holds $\mathbb{P}_{0}$-$a.s.$ that $m^{-1}\sum_{k=1}^{m}\mathbb{E}_{0}[Y_{k}]\to c$, as $m\to+\infty$.
Proof of Lemma 7.1.
It holds by the Chebyshev inequality that for every $\epsilon>0$,
$\sum_{m=1}^{+\infty}\mathbb{P}_{0}\left(\left|\frac{1}{m^{2}}\sum_{k=1}^{m^{2}}\left(Y_{k}-\mathbb{E}_{0}\left[Y_{k}\right]\right)\right|\geq\epsilon\right)=\sum_{m=1}^{+\infty}\frac{1}{\epsilon^{2}m^{4}}\sum_{k=1}^{m^{2}}{\rm Var}_{0}(Y_{k})\\ +\sum_{m=1}^{+\infty}\frac{1}{\epsilon^{2}m^{4}}\sum_{k_{1}\neq k_{2}}^{m^{2}}\mathbb{E}_{0}\left[\left(Y_{k_{1}}-\mathbb{E}_{0}\left[Y_{k_{1}}\right]\right)\left(Y_{k_{2}}-\mathbb{E}_{0}\left[Y_{k_{2}}\right]\right)\right]<+\infty,$
where we have applied asymptotic pairwise uncorrelation. Hence, the Borel-Cantelli lemma yields $\lim_{m\to+\infty}m^{-2}\sum_{k=1}^{m^{2}}(Y_{k}-\mathbb{E}_{0}[Y_{k}])=0,$ $\mathbb{P}_{0}$-$a.s.$ In order to de-square the number of summands $m^{2}$, let $m\in\mathbb{N}$ and let $n_{m}$ be the smallest index such that $m\in\{n_{m}^{2},\cdots,(n_{m}+1)^{2}\}$. Then, decompose the sum into two parts:
$\frac{1}{m}\sum_{k=1}^{m}\left(Y_{k}-\mathbb{E}_{0}\left[Y_{k}\right]\right)=\frac{1}{m}\sum_{k=1}^{(n_{m}+1)^{2}}\left(Y_{k}-\mathbb{E}_{0}\left[Y_{k}\right]\right)-\frac{1}{m}\sum_{k=m+1}^{(n_{m}+1)^{2}}\left(Y_{k}-\mathbb{E}_{0}\left[Y_{k}\right]\right),$
each of which vanishes $\mathbb{P}_{0}$-$a.s.$ as $m\to+\infty$, because $(n_{m}+1)^{2}\sim m$ as $m\to+\infty$. ∎
Proof of Theorem 4.6.
Thanks to $\inf_{t\in\mathbb{N}}v^{2}_{0}(t)>0$ (Assumption 4.1) and $v_{1}^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}1$ (Theorem 4.4), it suffices to show that $v_{2}^{2}(t)-v_{1}^{2}(t)\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}0$. Let $\{\xi_{k}\}_{k\in\mathbb{N}}$ and $\{\eta_{k}\}_{k\in\mathbb{N}}$ be sequences of random variables defined by $\xi_{k}:=X_{k}-\mu$ and $\eta_{k}:=X_{k}^{2}-\mathbb{E}_{k-1}[X_{k}^{2}]$, with which we obtain
$v_{2}^{2}(t)-v_{1}^{2}(t)=\frac{1}{|M(t)|}v_{2}^{2}(t)+\frac{2\mu}{|M(t)|}\sum_{k\in M(t)}\xi_{k}+\frac{1}{|M(t)|}\sum_{k\in M(t)}\eta_{k}+\mu^{2}-\mu^{2}(t),\quad t\in\mathbb{N}.$
It holds by Lemma 7.1 that the first two terms tend to vanish. The third term tends to zero due to the second condition given in Theorem 4.6. Finally, it holds by Theorem 2.4 and the continuous mapping theorem that the last two terms tend to cancel as $t\to+\infty$. ∎
Finally, we validate the asymptotic results of the three stopping rules (4.3), (4.7), and (4.11). We deal with the first one (4.7) separately from the remaining three, since the first one is deterministic whereas the other two are probabilistic along with the strictly positive inflation $a(t)$.
Proof of Proposition 4.3.
The first claim is trivial, since every element of $\{X_{k}\}_{k\in\mathbb{N}}$ is not degenerate due to $\inf_{t\in\mathbb{N}}v^{2}_{0}(t)>0$ (Assumption 4.1). Next, observe that $\mathbb{P}_{0}(|\mu(\tau_{\epsilon,\delta}^{0}-1)-\mu|>\epsilon)>\delta$ and $\mathbb{P}_{0}(|\mu(\tau_{\epsilon,\delta}^{0})-\mu|>\epsilon)\leq\delta$, by the definition (4.3). Along with $\tau_{\epsilon,\delta}^{0}\to+\infty$ as $\epsilon\to 0+$ and the martingale central limit theorem (4.2), those can be approximated as follows:
$\delta\lesssim 2(1-\Phi(\epsilon\sqrt{|M(\tau_{\epsilon,\delta}^{0}-1)|}/v(\tau_{\epsilon,\delta}^{0}-1))),$
$2(1-\Phi(\epsilon\sqrt{|M(\tau_{\epsilon,\delta}^{0})|}/v(\tau_{\epsilon,\delta}^{0})))\lesssim\delta,$
for sufficiently small $\epsilon$. Hence, we get
$(\Phi^{-1}(1-\delta/2))^{2}\lesssim\frac{\epsilon^{2}|M(\tau^{0}_{\epsilon,\delta})|}{v^{2}(\tau^{0}_{\epsilon,\delta})}=\frac{\epsilon^{2}|M(\tau^{0}_{\epsilon,\delta}-1)|}{v^{2}(\tau^{0}_{\epsilon,\delta}-1)}\frac{|M(\tau^{0}_{\epsilon,\delta})|}{|M(\tau^{0}_{\epsilon,\delta}-1)|}\frac{v^{2}(\tau^{0}_{\epsilon,\delta}-1)}{v^{2}(\tau^{0}_{\epsilon,\delta})}\\ \lesssim\Phi^{-2}(1-\delta/2)\frac{|M(\tau^{0}_{\epsilon,\delta})|}{|M(\tau^{0}_{\epsilon,\delta}-1)|}\frac{v^{2}(\tau^{0}_{\epsilon,\delta}-1)}{v^{2}(\tau^{0}_{\epsilon,\delta})}\to(\Phi^{-1}(1-\delta/2))^{2},$ (7.2)
as $\epsilon\to 0+$, due to $v^{2}(t+1)/|M(t+1)|\sim v^{2}_{0}(t)/|M(t)|$ as $t\to+\infty$ (Assumption 4.1). This yields Part (ii). Finally, for Part (iii), observe that
$\epsilon^{1/q}\tau_{\epsilon,\delta}^{0}$ $=\epsilon^{1/q}\mathop{\rm argmin}_{t\in\mathbb{N}}\left\{\mathbb{P}_{0}\left(\sqrt{|M(t)|}\frac{|\mu(t)-\mu|}{v_{0}(t)}>\sqrt{|M(t)|}\frac{\epsilon}{v_{0}(t)}\right)\leq\delta\right\}$
$\sim\mathop{\rm argmin}_{t\in(0,+\infty)}\left\{2\left(1-\Phi\left(\epsilon\sqrt{|M(\lfloor\epsilon^{-1/q}t\rfloor)|}/v(\lfloor\epsilon^{-1/q}t\rfloor)\right)\right)\leq\delta\right\},$
which yields the first result as $\epsilon\to 0+$, due to the martingale central limit theorem (4.2), as well as the convergence $\epsilon\sqrt{|M(\lfloor\epsilon^{-1/q}t\rfloor)|}/v(\lfloor\epsilon^{-1/q}t\rfloor)\to t^{q}/\sqrt{c}$ as $\epsilon\to 0+$. Finally, observe that
$\epsilon^{-1}(\mu(\tau_{\epsilon,\delta}^{0})-\mu)$ $\sim\frac{(\tau_{\epsilon,\delta}^{0})^{q}}{\sqrt{c}\Phi^{-1}(1-\delta/2)}(\mu(\tau_{\epsilon,\delta}^{0})-\mu)$
$\sim\frac{1}{\Phi^{-1}(1-\delta/2)}\sqrt{|M(\tau_{\epsilon,\delta}^{0})|}\frac{\mu(\tau_{\epsilon,\delta}^{0})-\mu}{v(\tau_{\epsilon,\delta}^{0})},$ (7.3)
which yields the desired weak convergence due to the martingale central limit theorem (4.2) and $\tau_{\epsilon,\delta}^{0}\to+\infty$. ∎
Proof of Proposition 4.5.
For Proposition 4.5, the expression (4.7) can be rewritten as $\tau^{a}_{\epsilon,\delta}=\min_{t\in\mathbb{N}}\{\epsilon\sqrt{|M(t)|}/(v(t)+a(t))\geq\Phi^{-1}(1-\delta/2)\}$, which yields $\tau^{a}_{\epsilon,\delta}\stackrel{{\scriptstyle a.s.}}{{\to}}+\infty$ as $\epsilon\to 0+$ or $\delta\to 0+$, due to $|M(t)|\to+\infty$ (Assumption 2.3), $v^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle a.s.}}{{\to}}1$ (Theorem 4.4) and $\inf_{t\in\mathbb{N}}v^{2}_{0}(t)>0$ (Assumption 4.1). With the aid of this expression, we get $\epsilon\sqrt{|M(\tau^{a}_{\epsilon,\delta})|}/(v(\tau^{a}_{\epsilon,\delta})+a(\tau^{a}_{\epsilon,\delta}))\geq\Phi^{-1}(1-\delta/2)$ at the time $m(\tau^{a}_{\epsilon,\delta})$, and $\epsilon\sqrt{|M(\tau^{a}_{\epsilon,\delta}-1)|}/(v(\tau^{a}_{\epsilon,\delta}-1)+a(\tau^{a}_{\epsilon,\delta}-1))<\Phi^{-1}(1-\delta/2)$ at the time $m(\tau^{a}_{\epsilon,\delta}-1)$, which together give, in a similar manner to (7.2),
$\Phi^{-1}(1-\delta/2)\leq\frac{\epsilon\sqrt{|M(\tau^{a}_{\epsilon,\delta})|}}{v(\tau^{a}_{\epsilon,\delta})+a(\tau^{a}_{\epsilon,\delta})}$
$\,=\frac{\epsilon\sqrt{|M(\tau^{a}_{\epsilon,\delta}-1)|}}{v(\tau^{a}_{\epsilon,\delta}-1)+a(\tau^{a}_{\epsilon,\delta}-1)}\frac{\sqrt{|M(\tau^{a}_{\epsilon,\delta})|}}{\sqrt{|M(\tau^{a}_{\epsilon,\delta}-1)|}}\frac{v(\tau^{a}_{\epsilon,\delta}-1)+a(\tau^{a}_{\epsilon,\delta}-1)}{v(\tau^{a}_{\epsilon,\delta})+a(\tau^{a}_{\epsilon,\delta})}$
$\,<\Phi^{-1}(1-\delta/2)\frac{\sqrt{|M(\tau^{a}_{\epsilon,\delta})|}}{\sqrt{|M(\tau^{a}_{\epsilon,\delta}-1)|}}\frac{v(\tau^{a}_{\epsilon,\delta}-1)+a(\tau^{a}_{\epsilon,\delta}-1)}{v(\tau^{a}_{\epsilon,\delta})+a(\tau^{a}_{\epsilon,\delta})}\stackrel{{\scriptstyle a.s.}}{{\to}}\Phi^{-1}(1-\delta/2),$ (7.4)
as $\epsilon\to 0+$, due to the first result $\tau^{a}_{\epsilon,\delta}\stackrel{{\scriptstyle a.s.}}{{\to}}+\infty$, as well as $v^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle a.s.}}{{\to}}1$ (Theorem 4.4), $v^{2}(t+1)/|M(t+1)|\sim v^{2}_{0}(t)/|M(t)|$ and $\inf_{t\in\mathbb{N}}v^{2}_{0}(t)>0$ (Assumption 4.1), and $\lim_{t\to+\infty}a(t)=0$. Hence, Part (ii) holds by the continuous mapping theorem since $\Phi$ is continuous on $\mathbb{R}$. With the additional condition, the first result of Part (iii) can be obtained by reformulating the equation (7.4). The second result of Part (iii) holds by the asymptotic equivalence (7.3) with the resampled empirical mean $\mu_{\star}$ and the stopping rule $\tau_{\epsilon,\delta}^{a}$, which yields the desired weak convergence due to the martingale central limit theorem (4.6) with the aid of $v^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle a.s.}}{{\to}}1$ (Theorem 4.4) and $\tau_{\epsilon,\delta}^{a}\stackrel{{\scriptstyle a.s.}}{{\to}}+\infty$ (Proposition 4.5 (i)), where the two asymptotic equivalences above hold almost surely, respectively, due to the first result and the additional condition of Part (iii). ∎
Supplementary materials: Stopping rules for Monte Carlo methods of martingale difference type
We provide supplementary materials for “Stopping rules for Monte Carlo methods of martingale difference type,” which we skipped in the main text to maintain the flow of the paper.
Appendix A Verification of technical assumptions and conditions
We here collect verifications of the standing assumptions (Assumptions 2.1, 2.3 and 4.1) and the technical conditions (Theorems 2.4, 4.2, 4.4 and 4.6) imposed along the way on the two examples examined in Section 5.
A.1 ARCH(1)
Recall that $X_{k}=(\beta+\alpha X_{k-1}^{2})^{1/2}V_{k}$ for all $k\in\mathbb{N}$ with $X_{0}=0$ and $\{V_{k}\}_{k\in\mathbb{N}}$ is a sequence of iid random variables with the $t$-distribution with $n$ degrees of freedom for some $n\in\{5,6,\cdots\}$, scaled in such a way to satisfy $\mathbb{E}_{0}[V_{1}^{2}]=1$, that is, the common probability density function is given by
$\frac{d}{dx}\mathbb{P}(V_{1}\leq x)=\frac{\Gamma((n+1)/2)}{\sqrt{(n-2)\pi}\Gamma(n/2)}\left(1+\frac{x^{2}}{n-2}\right)^{-(n+1)/2},\quad x\in\mathbb{R}.$ (A.1)
In particular, we have $\mathbb{E}_{0}[|V_{1}|^{4}]=3(n-2)/(n-4)$. It is trivial to derive that $\mathbb{E}_{k-1}[X_{k}]=(\beta+\alpha X_{k-1}^{2})^{1/2}\mathbb{E}_{k-1}[V_{k}]=0$ and $\mathbb{E}_{k-1}[X_{k}^{2}]=\beta+\alpha X_{k-1}^{2}$, due to $X_{k-1}\in\mathcal{F}_{k-1}$, $V_{k}\perp\mathcal{F}_{k-1}$, and $\mathbb{E}_{k-1}[V_{k}^{2}]={\rm Var}_{k-1}(V_{k})={\rm Var}_{0}(V_{k})=1$ for all $k\in\mathbb{N}$. Hence, we get $\mathbb{E}_{0}[X_{k}]=0$ (Assumption 2.1 with $\mu=0$) and ${\rm Var}_{0}(X_{k})=\mathbb{E}_{0}[X_{k}^{2}]=\beta+\alpha\mathbb{E}_{0}[X_{k-1}^{2}]=\cdots=\beta(1-\alpha^{k})/(1-\alpha)>0$, due to $\alpha\in(0,1)$, $\beta\in(0,+\infty)$ and $X_{0}=0$. Hence, the discrete-time stochastic process $\{X_{k}:\,k\in\mathbb{N}_{0}\}$ is asymptotically weakly stationary with mean $\mathbb{E}_{0}[X_{k}]=0$ and variance $\lim_{k\to+\infty}{\rm Var}_{0}(X_{k})=\beta/(1-\alpha)$, due to $\alpha\in(0,1)$. Hence, we get the convergence of the theoretical batch variance as
$v^{2}_{0}(t)=\frac{1}{|M(t)|}\sum_{k\in M(t)}{\rm Var}_{0}(X_{k})=\frac{\beta}{1-\alpha}\left(1-\frac{\alpha^{m(t-1)+1}-\alpha^{m(t)}}{2}\right)\to\frac{\beta}{1-\alpha},$
as $t\to+\infty$ due to $m(t)\to+\infty$. Hence, we get the assumptions $\inf_{t\in\mathbb{N}}v^{2}_{0}(t)>0$ and $\sup_{t\in\mathbb{N}}v^{2}_{0}(t)<+\infty$ (Assumption 4.1), and thus $\limsup_{t\to+\infty}\sum_{k\in M(t)}(k-m(t-1))^{-2}{\rm Var}_{0}(X_{k})<+\infty$ (Theorem 2.4). In addition to $|M(t)|\to+\infty$ (Assumption 2.3), we need to set the batch sizes $\{M(t)\}_{k\in\mathbb{N}}$ such that $|M(t+1)|\sim|M(t)|$ to ensure $v^{2}(t+1)/M(t+1)\sim v^{2}_{0}(t)/M(t)$ (Assumption 4.1).
One can show that all the relevant conditions hold true if the constant $\alpha$ and the degree of freedom $n$ satisfy $3\alpha^{2}(n-2)/(n-4)<1$. To verify the convergence of the conditional batch variance, observe first that for every $k\in\mathbb{N}$ and $m\in\{k+1,\cdots\}$,
$\mathbb{E}_{0}\left[\left(X_{k}^{2}-\mathbb{E}_{0}\left[X_{k}^{2}\right]\right)\left(X_{m}^{2}-\mathbb{E}_{0}\left[X_{m}^{2}\right]\right)\right]$ $=\mathbb{E}_{0}\left[\left(X_{k}^{2}-\mathbb{E}_{0}\left[X_{k}^{2}\right]\right)\mathbb{E}_{k}\left[X_{m}^{2}-\mathbb{E}_{0}\left[X_{m}^{2}\right]\right]\right]$
$=\mathbb{E}_{0}\left[\left(X_{k}^{2}-\mathbb{E}_{0}\left[X_{k}^{2}\right]\right)\left(\beta+\alpha\mathbb{E}_{k}\left[X_{m-1}^{2}\right]-\mathbb{E}_{0}\left[X_{m}^{2}\right]\right)\right]$
$=\cdots=\alpha\mathbb{E}_{0}\left[\left(X_{k}^{2}-\mathbb{E}_{0}\left[X_{k}^{2}\right]\right)\left(X_{m-1}^{2}-\mathbb{E}_{0}\left[X_{m-1}^{2}\right]\right)\right]$
$=\cdots=\alpha^{m-k}{\rm Var}_{0}(X_{k}^{2}),$
which tends to zero a lot faster than $m^{-2}$, due to $|\alpha|<1$, if $\sup_{k\in\mathbb{N}}{\rm Var}_{0}(X_{k}^{2})<+\infty$. To verify $\sup_{k\in\mathbb{N}}{\rm Var}_{0}(X_{k}^{2})<+\infty$, observe that for every $k\in\mathbb{N}$,
$\mathbb{E}_{0}\left[|X_{k}|^{4}\right]$ $=\mathbb{E}_{0}\left[(\beta+\alpha X_{k-1}^{2})^{2}\right]\mathbb{E}_{0}\left[|V_{k}|^{4}\right]$
$=\frac{3(n-2)}{n-4}\left(\beta^{2}+2\beta\alpha\mathbb{E}_{0}\left[X_{k-1}^{2}\right]+\alpha^{2}\mathbb{E}_{0}\left[|X_{k-1}|^{4}\right]\right)$
$\sim\frac{3(n-2)}{n-4}\left(\beta^{2}\frac{1+\alpha}{1-\alpha}+\alpha^{2}\mathbb{E}_{0}\left[|X_{k-1}|^{4}\right]\right).$
If $3\alpha^{2}(n-2)/(n-4)<1$, then we get
$\mathbb{E}_{0}\left[|X_{k}|^{4}\right]\to\frac{3\beta^{2}(1+\alpha)(n-2)}{(1-\alpha)(n-4-3\alpha^{2}(n-2))},$
${\rm Var}_{0}\left(X_{k}^{2}\right)\to\frac{3\beta^{2}(1+\alpha)(n-2)}{(1-\alpha)(n-4-3\alpha^{2}(n-2))}-\frac{\beta^{2}}{(1-\alpha)^{2}},$
as $k\to+\infty$. Hence, by Lemma 7.1, we get $\lim_{t\to+\infty}|M(t)|^{-1}\sum_{k\in M(t)}X_{k-1}^{2}=\lim_{k\to+\infty}\mathbb{E}_{0}[X_{k}^{2}]=\beta/(1-\alpha)$. That is, with $\mu=0$, we get
$v_{1}^{2}(t)$ $=\frac{1}{|M(t)|}\sum_{k\in M(t)}\mathbb{E}_{k-1}\left[X_{k}^{2}\right]-\mu^{2}=\frac{1}{|M(t)|}\sum_{k\in M(t)}\left(\beta+\alpha X_{k-1}^{2}\right)$
$=\beta+\frac{\alpha}{|M(t)|}\sum_{k\in M(t)}X_{k-1}^{2}\to\frac{\beta}{1-\alpha},$
$\mathbb{P}_{0}$-$a.s.$, as $t\to+\infty$. With $v^{2}_{0}(t)\to\beta/(1-\alpha)$, the condition $v_{1}^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}1$ (Theorem 4.4) holds true. In addition, the convergence of the sequence $\{\mathbb{E}_{0}[|X_{k}|^{4}]\}_{k\in\mathbb{N}}$ above implies the condition $\limsup_{k\to+\infty}\mathbb{E}_{0}[|X_{k}|^{4}]<+\infty$ (Theorem 4.6). Now, with $c_{n}$ being a constant depending on $n$ changing its values from line to line, due to
$\mathbb{E}_{k-1}\left[V_{k}^{2}\mathbbm{1}(|X_{k}|>\epsilon\sqrt{|M(t)|})\right]$ $=c_{n}\int_{x^{2}>\frac{\epsilon^{2}|M(t)|}{\beta+\alpha X_{k-1}^{2}}}x^{2}\left(1+\frac{x^{2}}{n-2}\right)^{-(n+1)/2}dx$
$\sim c_{n}\left(\frac{\epsilon^{2}|M(t)|}{\beta+\alpha X_{k-1}^{2}}\right)^{1-n/2},$
as $t\to+\infty$ with the aid of the probability density function (A.1), we get
$\frac{1}{|M(t)|}\sum_{k\in M(t)}\mathbb{E}_{k-1}\left[X_{k}^{2}\mathbbm{1}(|X_{k}|>\epsilon\sqrt{|M(t)|})\right]$
$\qquad=\frac{1}{|M(t)|}\sum_{k\in M(t)}\left(\beta+\alpha X_{k-1}^{2}\right)\mathbb{E}_{k-1}\left[V_{k}^{2}\mathbbm{1}(|X_{k}|>\epsilon\sqrt{|M(t)|})\right]$
$\qquad\sim c_{n}\epsilon^{2-n}\sum_{k\in M(t)}\left(\frac{\beta+\alpha X_{k-1}^{2}}{|M(t)|}\right)^{n/2},$
which tends to 0, $\mathbb{P}_{0}$-$a.s.$, since $\sum_{k\in M(t)}(\beta+\alpha X_{k-1}^{2})/|M(t)|$ is convergent to a strictly positive constant and $n\in\{5,6,\cdots\}$. This ensures the Lindeberg condition of Theorem 4.2. Next, to ensure the second condition of Theorem 4.6, observe that
$\frac{1}{|M(t)|}\sum_{k\in M(t)}\left(X_{k}^{2}-\mathbb{E}_{k-1}\left[X_{k}^{2}\right]\right)=\frac{1}{|M(t)|}\sum_{k\in M(t)}X_{k}^{2}-\beta-\alpha\frac{1}{|M(t)|}\sum_{k\in M(t)}X_{k-1}^{2}\to 0,$
$\mathbb{P}_{0}$-$a.s.$, due to $|M(t)|^{-1}\sum_{k\in M(t)}X_{k-1}^{2}\to\beta/(1-\alpha)$.
A.2 Adaptive control variates
The theoretical and conditional batch variances are given, respectively, by
$v^{2}_{0}(t)$ $=\frac{1}{|M(t)|}\sum_{k\in M(t)}\mathbb{E}_{0}\left[X_{k}^{2}\right]-\mu^{2}$
$=V(0)+2\left\langle\mathbb{E}_{0}[{\bm{\theta}}(t-1)],\mathbb{E}_{0}\left[\Psi(U_{1})(U_{1}-\mathbbm{1}_{d}/2)\right]\right\rangle+\frac{1}{12}\mathbb{E}_{0}\left[\|{\bm{\theta}}(t-1)\|^{2}\right]-\mu^{2},$
and
$v_{1}^{2}(t)$ $=\frac{1}{|M(t)|}\sum_{k\in M(t)}\mathbb{E}_{k-1}\left[X_{k}^{2}\right]-\mu^{2}$
$=V(0)+2\left\langle{\bm{\theta}}(t-1),\mathbb{E}_{0}\left[\Psi(U_{1})(U_{1}-\mathbbm{1}_{d}/2)\right]\right\rangle+\frac{1}{12}\left\|{\bm{\theta}}(t-1)\right\|^{2}-\mu^{2},$
for all $t\in\mathbb{N}$ and $k\in M(t)$. By restricting the parameter domain to a sufficiently large yet compact one $\Theta(\subset\mathbb{R}^{d})$ (which, at least, contains the initial state $0_{d}$ and the optimizer $\bm{\theta}$$^{*}$), it holds that $v^{2}_{0}(t)\to V({\bm{\theta}}^{*})-\mu^{2}$ as well as $v_{1}^{2}(t)\to V({\bm{\theta}}^{*})-\mu^{2}$ by the bounded convergence theorem since the search domain $\Theta$ is compact with the aid of ${\bm{\theta}}(t)\to{\bm{\theta}}^{*}$, $\mathbb{P}_{0}$-$a.s.$ Hence, we get $\inf_{t\in\mathbb{N}}v^{2}_{0}(t)>0$ and $\sup_{t\in\mathbb{N}}v^{2}_{0}(t)<+\infty$ (Assumption 4.1), as well as $\sum_{k\in M(t)}(k-m(t-1))^{-2}{\rm Var}_{0}(X_{k})\sim\mathbb{E}_{0}[V({\bm{\theta}}^{*})]/|M(t)|\to 0$ (Theorem 2.4). One needs to set the batch sizes $\{M(t)\}_{t\in\mathbb{N}}$ satisfying $|M(t)|\to+\infty$ (Assumption 2.3) and $|M(t+1)|\sim|M(t)|$, to ensure $v^{2}(t+1)/|M(t+1)|\sim v^{2}_{0}(t)/|M(t)|$ (Assumption 2.3). Clearly, the condition $v_{1}^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}1$ (Theorem 4.4) is satisfied.
Due to the quadratic integrability condition $\int_{(0,1)^{d}}|\Psi({\bf u})|^{4}d{\bf u}<+\infty$, it immediately holds that $\limsup_{k\to+\infty}\mathbb{E}_{0}[|X_{k}|^{4}]<+\infty$ (Theorem 4.6) and that $\mathbb{E}_{k-1}[|X_{k}-\mu|^{4}]$ is uniformly bounded from both below and above, since the search domain $\Theta$ and the uniform random vectors $\{U_{k}\}_{k\in\mathbb{N}}$ are bounded. Hence, the Lindeberg condition (Theorem 4.2) holds by the convergence $|M(t)|^{-1}\sum_{k\in M(t)}\mathbb{E}_{k-1}[|X_{k}-\mu|^{4}]^{1/2}(\mathbb{P}_{k-1}(|X_{k}-\mu|>\epsilon\sqrt{|M(t)|}))^{1/2}\to 0$ (Theorem 4.2). Finally, it is straightforward to show that $|M(t)|^{-1}\sum_{k\in M(t)}(X_{k}^{2}-\mathbb{E}_{k-1}[X_{k}^{2}])\to 0$, $\mathbb{P}_{0}$-$a.s.$ (Theorem 4.6), due to ${\bm{\theta}}(t-1)\in\mathcal{F}_{k-1}$ and $U_{k}\perp\mathcal{F}_{k-1}$ for all $t\in\mathbb{N}$ and $k\in M(t)$.
Appendix B Comparisons with relevant stopping rules
We here contrast the proposed framework with two relevant stopping rules in order to put the proposed ones into perspective. In what follows, we denote by $\mathbb{D}((0,+\infty);\mathbb{R})$ the space of right-continuous functions in $\mathbb{R}$ with left limits defined on the open interval $(0,+\infty)$.
B.1 Stopping rules in asymptotic regimes
We have justified our formulations in an asymptotic manner along the way (Propositions 4.3 and 4.5) in a similar spirit to the existing framework in the asymptotic regime [12]. A key assumption can be rephrased in our context that the batch variances are convergent. That is, given that a limiting unit-time variance exits ($\lim_{t\to+\infty}v^{2}_{0}(t)(=:c)$), the strong consistency holds ($v_{1}^{2}(t)\stackrel{{\scriptstyle a.s.}}{{\to}}c$, or $v_{2}^{2}(t)\stackrel{{\scriptstyle a.s.}}{{\to}}c$), or the functional weak law of large numbers holds ($\{s^{2}(\lfloor t/\epsilon\rfloor):\,t\in(0,+\infty)\}\stackrel{{\scriptstyle\mathcal{L}}}{{\to}}c$, or $\{\sigma^{2}(\lfloor t/\epsilon\rfloor):\,t\in(0,+\infty)\}\stackrel{{\scriptstyle\mathcal{L}}}{{\to}}c$ in $\mathbb{D}((0,+\infty);\mathbb{R})$). Clearly, all those conditions are stronger than ours (Assumption 4.1 and the additional conditions in Proposition 4.5). Another key condition is the existence of a limiting continuous process of the scaled-centered estimation process weakly in $\mathbb{D}((0,+\infty);\mathbb{R}))$ in conjunction with the limiting unit-time variance. To verify this condition in our context, consider the (non-batched) empirical mean $\mu_{n}:=n^{-1}\sum_{k=1}^{n}X_{k}$, instead of the batched one (2.2), so that $\mu_{n}\to\mu$ almost surely by Theorem 2.4. For $\epsilon\in(0,+\infty)$, define stochastic process $\{\mathcal{Y}_{\epsilon}(t):\,t\in(0,+\infty)\}$ by
$\mathcal{Y}_{\epsilon}(t):=\epsilon^{-1/2}(\mu_{\lfloor t/\epsilon\rfloor}-\mu)=\epsilon^{-1/2}\frac{1}{\lfloor t/\epsilon\rfloor}\sum_{k=1}^{\lfloor t/\epsilon\rfloor}(X_{k}-\mathbb{E}_{k-1}[X_{k}]),\quad t\in(0,+\infty),$
which converges weakly to $\{W_{t}/t:\,t\in(0,+\infty)\}$ in $\mathbb{D}((0,+\infty);\mathbb{R})$, where $\{W_{t}:\,t\geq 0\}$ is a centered Brownian motion with ${\rm Var}_{0}(W_{1})=c(=\lim_{t\to+\infty}v^{2}_{0}(t))$. To derive this weak convergence, consider $\mathcal{X}_{\epsilon}(t):=\epsilon\lfloor t/\epsilon\rfloor\mathcal{Y}_{\epsilon}(t)=\epsilon^{1/2}\sum_{k=1}^{\lfloor t/\epsilon\rfloor}(X_{k}-\mathbb{E}_{k-1}[X_{k}])$. The stochastic process $\{\mathcal{X}_{\epsilon}(t):\,t\in(0,+\infty)\}$ is centered with uncorrelated and asymptotically second-order stationary increments, due to $\mathbb{E}_{0}[\mathcal{X}_{\epsilon}(t_{1})]=0$, ${\rm Cov}_{0}(\mathcal{X}_{\epsilon}(t_{2})-\mathcal{X}_{\epsilon}(t_{1}),\mathcal{X}_{\epsilon}(t_{4})-\mathcal{X}_{\epsilon}(t_{3}))=0$ and
${\rm Var}_{0}(\mathcal{X}_{\epsilon}(t_{2})-\mathcal{X}_{\epsilon}(t_{1}))$ $=\epsilon(\lfloor t_{2}/\epsilon\rfloor-\lfloor t_{1}/\epsilon\rfloor)\frac{1}{\lfloor t_{2}/\epsilon\rfloor-\lfloor t_{1}/\epsilon\rfloor}\sum_{k=\lfloor t_{1}/\epsilon\rfloor+1}^{\lfloor t_{2}/\epsilon\rfloor}{\rm Var}_{0}(X_{k})$
$\to(t_{2}-t_{1})c,$
for all $0\leq t_{1}\leq t_{2}\leq t_{3}\leq t_{4}$. To further show that the limiting process is a Brownian motion, let $\{\epsilon_{n}\}_{n\in\mathbb{N}}$ and $\{h_{n}\}_{n\in\mathbb{N}}$ be sequences of positive constants with $\epsilon_{n}\to 0$ and $h_{n}\to 0$ as $n\to+\infty$ and let $\tau_{n}$ be a stopping time with respect to the stochastic process $\{\mathcal{X}_{\epsilon_{n}}(t):\,t\geq 0\}$ for all $n\in\mathbb{N}$, that is, $\{\tau_{n}\leq t\}\in\sigma(\{\mathcal{X}_{\epsilon_{n}}(s):\,s\in[0,t]\})$. Due to the martingale central limit theorems (Theorems 4.2, 4.4 and 4.6), the limiting process is Gaussian and $\{\mathcal{X}_{\epsilon_{n}}(t):\,t\geq 0\}$ converges weakly in $\mathbb{D}([0,+\infty);\mathbb{R})$ to a centered Brownian motion, due to tightness $\mathbb{P}_{0}(|\mathcal{X}_{\epsilon_{n}}(\tau_{n}+h_{n})-\mathcal{X}_{\epsilon_{n}}(\tau_{n})|>m)\lesssim m^{-2}h_{n}\lim_{t\to+\infty}v^{2}_{0}(t)\to 0$ as $n\to+\infty$ for all $m>0$. (See, for instance, [20, Chapter 23].)
B.2 Relative-precision stopping rules
[…]
$\delta<\mathbb{P}_{m(\tau)}\left(|\mu_{\star}(\tau)-\mu|>\epsilon|\mu|\right)$ $=\mathbb{P}_{m(\tau)}\left(\sqrt{|M(\tau)|}\frac{|\mu_{\star}(\tau)-\mu|}{v(\tau)}>\sqrt{|M(\tau)|}\frac{\epsilon|\mu|}{v(\tau)}\right)$
$\sim 2\left(1-\Phi(\epsilon\sqrt{|M(\tau)|}|\mu|/v(\tau))\right),$
where we have employed the key fact that the resampled batch $\{X_{k}^{\star}\}_{k\in M(\tau)}$ is conditionally independent of the stopping batch $\tau$ on the stopping-time $\sigma$-field $\mathcal{F}_{m(\tau)}$ for the martingale central limit theorem (Theorem 4.2). The crucial issue here lies in how to treat the unknown mean $\mu$ in the rightmost term, which is exactly what one is looking for. On the one hand, it cannot be replaced with the original empirical mean $\mu(\tau)$ because it is not clear whether $\mu(\tau)$ approximates $\mu$. We stress, on the other hand, that $\mu$ there cannot be replaced with the resampled empirical mean $\mu_{\star}(\tau)$ either despite it approximates $\mu$ on the stopping-time $\sigma$-field $\mathcal{F}_{m(\tau)}$, since then the stopping batch $\tau$ would be determined after generating the resampled one $\{X_{k}^{\star}\}_{k\in M(\tau)}$ in reverse order.
Before closing this section, let us note, as already outlined along Assumption 2.1, that a relative-precision stopping rule was developed in [9] upon a similar assumption to (2.1), but along with a boundedness condition on the underlying randomness (specifically, $X_{n}\in[0,1]$, $a.s.$), which plays an essential role for deriving bounds for the error probability upon which the stopping rule is built.